### v2

# On Mirzakhani’s twist torus conjecture

Abstract.
We address a conjecture of Mirzakhani about the statistical behavior of certain expanding families of “twist tori” in the moduli space of hyperbolic surfaces, showing that they equidistribute to a certain Lebesgue-class measure along almost all sequences. We also identify a number of other expanding families of twist tori whose limiting distributions are mutually singular to Lebesgue.
1. Introduction
Let $\mathcal{M}_{g}$ be the moduli space of hyperbolic structures on a closed, oriented surface $S$ of genus $g\geq 2$. Let $\gamma$ be a pants decomposition of $S$, i.e., a maximal collection of non-isotopic, pairwise disjoint, essential, simple closed curves. For any $L>0$, there is a twist torus $\mathbb{T}_{\gamma}(L)\subset\mathcal{M}_{g}$ consisting of those hyperbolic surfaces glued together from pairs of pants with all boundary components of length $L$ along a pants decomposition of topological type $\gamma$. The gluing/twist parameters take values in a torus $(\mathbb{R}/L\mathbb{Z})^{3g-3}$, as one full twist about any curve in $\gamma$ yields an isometric hyperbolic surface. The twist torus $\mathbb{T}_{\gamma}(L)$ is an immersed finite quotient [^1] As $\gamma$ may have a non-trivial finite symmetry group. of $(\mathbb{R}/L\mathbb{Z})^{3g-3}$, which we equip with the pushforward $\tau_{\gamma}(L)$ of the Lebesgue measure on $(\mathbb{R}/L\mathbb{Z})^{3g-3}$. See §3 for a more detailed discussion.
Mirzakhani conjectured [^2] See [Wri20, Problem 13.2]. that as $L\to\infty$, these tori equidistribute in $\mathcal{M}_{g}$ to a certain measure in the class of Lebesgue. See §2 for definitions of the function $B(X)$, the constant $b_{g}$, and the measure $\operatorname{vol}_{\operatorname{WP}}$.
Conjecture 1.1 (Mirzakhani’s twist torus conjecture).
Let $\gamma$ be any pants decomposition. Then
$\lim_{L\to\infty}\frac{\tau_{\gamma}(L)}{L^{3g-3}}=\frac{B(X)}{b_{g}}d\operatorname{vol}_{\operatorname{WP}}(X)$
with respect to the weak-$\ast$ topology on measures on $\mathcal{M}_{g}$.
This should be compared with a classical result regarding the equidistribution of expanding closed horocycles in the unit tangent bundle $\mathsf{SL}_{2}\mathbb{Z}\backslash\mathsf{SL}_{2}\mathbb{R}=T^{1}Y$ of the modular curve $Y$. Denote by $\Lambda(L)$ the Lebesgue measure on the (unique) closed orbit of the unstable horocycle flow of period $L$. Then $\Lambda(L)/L$ converges weak-* as $L\to\infty$ to the Liouville measure on $T^{1}Y$, normalized to have unit volume [Sar81]; see [EM93, §II] for a proof using ideas introduced in Margulis’s thesis [Mar70, Mar04] and mixing of the geodesic flow [Hed39]. Using the projection $p:T^{1}Y\to Y$, one obtains an equidistribution result of the form
$\lim_{L\to\infty}\frac{p_{*}\Lambda(L)}{L}=\frac{2\pi}{\operatorname{vol}(T^{1}Y)}d\operatorname{Area}_{Y},$
where $d\operatorname{Area}_{Y}$ is the hyperbolic area measure on $Y$.
Conjecture 1.1 also has a refinement in terms of the equidistribution of expanding orbits of a flow. Using geodesic length functions, Thurston identified the unit cotangent bundle of $\mathcal{M}_{g}$ with $\mathcal{P}^{1}\mathcal{M}_{g}$, the bundle of unit length measured geodesic laminations over $\mathcal{M}_{g}$ (see §2). This bundle carries the unipotent-like earthquake flow: given a hyperbolic surface $X$ and a measured lamination $\lambda$, the earthquake flow acts by cutting open $X$ along $\lambda$, shearing along $\lambda$ according to its measure, and then regluing. There is an earthquake flow–invariant ergodic measure $\mu_{\operatorname{Mirz}}$ on $\mathcal{P}^{1}\mathcal{M}_{g}$ in the class of Lebesgue, which satisfies
$dp_{*}\mu_{\operatorname{Mirz}}(X)={B(X)}~{}d\operatorname{vol}_{\operatorname{WP}}(X),$
where $p:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{M}_{g}$ forgets the measured lamination. Its mass is the constant $b_{g}$. Recent results have tied the ergodic theory of the earthquake flow with respect to $\mu_{\operatorname{Mirz}}$ to questions about counting hyperbolic geodesics [Mir07, AH20, Liu20, AHC22].
Twist tori lift to $\mathcal{P}^{1}\mathcal{M}_{g}$ by keeping track of the multicurve $\gamma$ weighted by $(L\cdot\#\gamma)^{-1}$. Evidently, these lifted twist tori are invariant under the earthquake flow. Conjecture 1.1 would therefore follow from the stronger statement that the normalized Lebesgue measures on lifted twist tori equidistribute to $\mu_{\operatorname{Mirz}}/b_{g}$ in $\mathcal{P}^{1}\mathcal{M}_{g}$ as $L\to\infty$.
In this paper, we address this stronger version as well as the following generalization. Fix a pants decomposition $\gamma=\gamma_{1}\cup\ldots\cup\gamma_{3g-3}$, a length vector $\bm{\ell}\in\mathbb{R}_{>0}^{3g-3}$, and a weight vector $\mathbf{h}\in\mathbb{R}^{3g-3}_{>0}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. Denote by $\mathbf{h}\gamma$ the measured lamination $h_{1}\gamma_{1}\cup\ldots\cup h_{3g-3}\gamma_{3g-3}$. Setting the length of each $\gamma_{i}$ equal to $\ell_{i}$ and taking arbitrary twists yields an immersion of a (finite quotient of a) torus into $\mathcal{M}_{g}$ as before, and we can lift this this to a map
$(\mathbb{R}/\ell_{1}\mathbb{Z})\times\ldots\times(\mathbb{R}/\ell_{3g-3}\mathbb{Z})\looparrowright\mathcal{P}^{1}\mathcal{M}_{g}$
by specifying the lamination coordinate to be $\mathbf{h}\gamma$. Let $\mathcal{P}\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})\subset\mathcal{P}^{1}\mathcal{M}_{g}$ denote the image of this map and let $\mu_{\gamma}(\bm{\ell},\mathbf{h})$ be the pushforward of Lebesgue measure, normalized to have unit mass. We observe that for almost every choice of weights $\mathbf{h}$, the measure $\mu_{\gamma}(\bm{\ell},\mathbf{h})$ is ergodic for the (translation) action of the earthquake flow on $\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})$. Again, see §3 for the precise constructions.
Say that a pair of pants has property ()($\nabla$)( ∇ ) if one of its boundary lengths is the sum of the other two.
Theorem 1.2.
Fix any pants decomposition $\gamma$, any >$\bm{\ell}\in\mathbb{R}^{3g-3}_{>0}$, and any $\mathbf{h}\in\mathbb{R}^{3g-3}_{>0}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. Then there is a set $Z\subset\mathbb{R}$ of zero density such that
$\lim_{\begin{subarray}{c}t\to\infty\\ t\notin Z\end{subarray}}\mu_{\gamma}(e^{t}\bm{\ell},e^{-t}\mathbf{h})=\mu_{\infty}$
where $\mu_{\infty}$ is an earthquake-flow invariant ergodic probability measure on $\mathcal{P}^{1}\mathcal{M}_{g}$. Moreover,
  1. (1)
    if $g\geq 3$ and no component of $S\setminus\gamma$ has ()($\nabla$)( ∇ ), then $\mu_{\infty}$ is $\mu_{\operatorname{Mirz}}/b_{g}$.
  2. (2)
    if some component of $S\setminus\gamma$ has ()($\nabla$)( ∇ ), then $\mu_{\infty}$ is not equal (hence singular) to $\mu_{\operatorname{Mirz}}/b_{g}$.
In particular, Conjecture 1.1 is true outside a set of times $e^{Z}$ where $Z$ has zero density.
Remark 1.3.
For $g=2$, statement (1) of Theorem 1.2 would follow, using our techniques, from [AW23, Conjecture 1.2] about “almost high rank” invariant submanifolds; see §§45.
The following averaged version of Conjecture 1.1 is an immediate consequence (just push down to $\mathcal{M}_{g}$):
[…]
The limiting measure $\mu_{\infty}$ always has more structure than just earthquake-flow invariance. In [CF24b], the authors extended the earthquake flow to a (measurable) action by “stretchquakes.” There and in [CF24a], the authors showed that stretchquake-invariant probabilty measures on $\mathcal{P}^{1}\mathcal{M}_{g}$ correspond to Lebesgue-class measures on affine $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant subvarieties of strata of holomorphic quadratic differentials (§4.2). The measure $\mu_{\infty}$ is always stretchquake-invariant, and most of the time, we can identify it as corresponding to the Masur–Smillie–Veech measure on a component of a stratum of quadratic differentials with some zeros of order 2; see Corollary 5.12. We emphasize that Mirzakhani’s conjecture only references $\mu_{\operatorname{Mirz}}$; it seems to have been unexpected that families of twist tori could equidistribute to anything else.
Figure 1. Hyperbolic and flat twist tori. These tori are mapped to each other under $\mathcal{O}$.
The strategy of our proof is to exploit a deep connection between the earthquake and horocycle flows discovered by Mirzakhani [Mir08] and further developed by the authors in [CF24b, CF24a]. Briefly, there is a Borel isomorphism $\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ that takes the earthquake flow to the horocycle flow on the moduli space $\mathcal{Q}^{1}\mathcal{M}_{g}$ of unit area quadratic differentials in a time-preserving way. While $\mathcal{O}$ is not continuous, it is continuous along the leaves of a certain foliation and each $\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})$ is contained in such a leaf. This allows us to map the twist torus via $\mathcal{O}$ to get a torus in $\mathcal{Q}^{1}\mathcal{M}_{g}$ consisting of flat surfaces glued together from horizontal cylinders corresponding to the components of $\gamma$ with heights $\mathbf{h}$ and lengths $\bm{\ell}$; see Figure 1. Once in this setting, we can use powerful results from Teichmüller dynamics to understand the limiting distribution of expanding twist tori in $\mathcal{Q}^{1}\mathcal{M}_{g}$.
The driving engine of our proof is then the following theorem, which allows us to pull back equidistribution statements from $\mathcal{Q}^{1}\mathcal{M}_{g}$ to $\mathcal{P}^{1}\mathcal{M}_{g}$. This is a special case of [CF24a, Theorem B].
Theorem 1.4.
Suppose that $\nu_{n}\to\nu$ is a convergent sequence of probability measures on a stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$, and suppose that $\nu$ is a $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$–invariant probability measure. Then $\mathcal{O}^{*}\nu_{n}\to\mathcal{O}^{*}\nu$ on $\mathcal{P}^{1}\mathcal{M}_{g}$.
Thus, having proved an equidistribution statement for expanding twist tori in $\mathcal{Q}^{1}\mathcal{M}_{g}$, we can simply pull it back to deduce Theorem 1.2. This scheme also works to show the equidistribution of many different families of expanding twist tori where $\gamma$ is an arbitrary multicurve, and can be used to show equidistribution to many different limiting distributions. See Theorems 5.13, 5.14, and 5.16.
1.1. Context: unipotent flows on moduli spaces
There are surprising similarities between the dynamics of groups generated by unipotents on homogeneous spaces and the action of $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$ on strata of holomorphic differentials. By way of [Mir08] and [CF24b, CF24a], parts of this analogy can now be extended to the action (of the group $P$ of upper triangular matrices in $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$) by stretchquakes on $\mathcal{P}^{1}\mathcal{M}_{g}$.
A subgroup $U$ of a connected reductive Lie group $G$ is horospherical if there is a $g\in G$ such that
$U=\{u\in G:g^{n}ug^{-n}\to id,~{}n\to\infty\}.$
For a lattice $\Gamma\leq G$ and horospherical subgroup $U\leq G$, the $U$-invariant ergodic probability measures and $U$-orbit closures in $\Gamma\backslash G$ were classified by Dani [Dan86]: they are homogeneous. In her landmark results, Ratner proved that such rigidity phenomena hold for arbitrary connected subgroups $H\leq G$ generated by unipotents [Rat91a, Rat91b]: every $H$-invariant ergodic probability measure and $H$-orbit closure in $\Gamma\backslash G$ is homogeneous.
Despite the seemingly insurmountable difficulty that strata of holomorphic quadratic differentials on higher genus surfaces are completely inhomogeneous, Eskin and Mirzakhani showed that $P$-invariant probability measures are extremely well behaved [EM18], and with Mohammadi they proved that the same is true for orbit closures [EMM15] (see §4.2 and Theorem 4.2 for a summary of their results).
Leveraging the ergodicity of the earthquake flow, Mirzakhani showed that the analogues of expanding horospheres ($L$-level sets of the hyperbolic length of a simple closed curve) equidistribute to $\pi_{*}\mu_{\operatorname{Mirz}}$ on $\mathcal{M}_{g}$ as $L\rightarrow\infty$ [Mir07]. These level sets can be lifted to $\mathcal{P}^{1}\mathcal{M}_{g}$ in a natural way by keeping track of the curve, as in §3.1. There are finite, earthquake flow–invariant measures on these lifted level sets that are uniformly expanded/contracted in forward/backward time by the “dilation flow,” the diagonal part of the stretchquake action (see the end of §2.3). The lifted level sets have the maximum dimension with this property; see also Theorems 2.1 and 2.3 that explain that the conjugacy $\mathcal{O}$ maps lifted level sets to leaves of the strong unstable foliation for the Teichmüller geodesic flow. These observations justify Mirzakhani’s use of the “horospherical” terminology in [Mir07] (compare [Wri20, §13.3]).
Twist tori are only $(3g-3)$-dimensional inside of the $(6g-7)$-dimensional expanding horospheres of the previous paragraph; as such, they should not be considered horospherical. The much more refined results of Eskin–Mirzkhani–Mohammadi, analogous to Ratner’s Theorems, are vital ingredients in our proof of Theorem 1.2 and the even more delicate examples presented in §5.4. Compare Figure 2.
Mirzakhani’s equidistribution results for expanding horospheres were generalized by both Arana-Herrera [AH21] (who clarified a technical point in [Mir07]) and Liu [Liu20] to different types of length functions associated to multicurves and by Arana-Herrera and the first author [AHC22] to the case where one puts restrictions on the geometry of the complementary subsurface.
Our techniques recover and generalize these results. Indeed, let $\gamma=\gamma_{1}\cup\ldots\cup\gamma_{k}$ be any multicurve and $\mathbf{h}\in\mathbb{R}^{k}_{>0}$ any weight vector. Consider the $L$-level set of the hyperbolic length function for $\mathbf{h}\gamma:=h_{1}\gamma_{1}\cup\ldots\cup h_{k}\gamma_{k}$ in the Teichmüller space $\mathcal{T}_{g}$; this admits a lift to the unit cotangent bundle $\mathcal{P}^{1}\mathcal{T}_{g}$ by setting the second coordinate equal to $\mathbf{h}\gamma/L$. Any probability measure supported on such a lifted level set that is absolutely continuous with respect to the Lebesgue class with continuous density and such that its restriction to each earthquake flow line is the restriction of Lebesgue to an interval is said to be horospherical. These level sets can also be stratified by the combinatorial type of the ribbon graph spine for $X\setminus\gamma$ (see §2 and §§3.23.3); any measure satisfying the properties above but on a substratum is called a horospherical slice measure. We use the same terms to denote the pushforwards of such measures down to $\mathcal{P}^{1}\mathcal{M}_{g}$. See Section 3.3 and Figure 2.
Figure 2. Projections to the length part of Fenchel–Nielsen coordinates. From left to right: the horospheres considered in [Mir07], the horospheres considered in [AHC22], the twist tori of Conjecture 1.1, and a horospherical slice measure on a lower-dimensional stratum.
As the dilation flow $\left(\operatorname{Dil}_{t}\right)$ normalizes the earthquake flow, pushforwards of horospherical (slice) measures by the dilation flow are themselves horospherical (slice) measures. Applying a result of Forni [For21, Theorem 1.6] (see also [EMM22, Proposition 3.2]), we arrive at the following:
Theorem 1.5.
Let $\sigma$ be any horospherical (slice) measure on $\mathcal{P}^{1}\mathcal{M}_{g}$. Then
$\lim_{t\to\infty}\left(\operatorname{Dil}_{t}\right)_{*}\sigma=\mu_{\infty}$
where $\mu_{\infty}$ is an earthquake-flow invariant ergodic probability measure on $\mathcal{P}^{1}\mathcal{M}_{g}$. Moreover,
  1. (1)
    If $\sigma$ is horospherical or supported on a top-dimensional stratum, then $\mu_{\infty}=\mu_{\operatorname{Mirz}}/b_{g}$.
  2. (2)
    If $\sigma$ is supported on a lower-dimensional stratum, $\mu_{\infty}$ is not equal (hence singular) to $\mu_{\operatorname{Mirz}}$.
In the second case, the limiting measure is the pushforward via $\mathcal{O}^{-1}$ of the Masur–Smillie–Veech measure on the stratum component $\mathcal{Q}$ of $\mathcal{Q}^{1}\mathcal{M}_{g}$ corresponding to the horospherical slice; see §4.3 for the proof. The generic pair $(X,\lambda)$ with respect to one of these measures is a hyperbolic surface equipped with a measured lamination that cuts it into a union of regular ideal polygons [CF24b, Corollary 1.2].
Acknowledgments
We are very grateful to Francisco Arana-Herrera for an inspiring discussion at the beginning of this project and to Paul Apisa for explaining to us how to deduce Corollary 5.4 from his work with Alex Wright. AC was supported by NSF grants DMS-2005328 and DMS-2202703. JF was supported by NSF grant DMS-2005328 and DFG grants 427903332 and 281071066 – TRR 191.
2. Flows on moduli spaces
In this section, we discuss basic dynamical features of the spaces $\mathcal{P}^{1}\mathcal{M}_{g}$ and $\mathcal{Q}^{1}\mathcal{M}_{g}$ and describe a Borel measurable bijection $\mathcal{O}:\mathcal{PM}_{g}\to\mathcal{QM}_{g}$ conjugating their dynamics. Throughout, $S$ is a closed, oriented surface of genus $g\geq 2$ and $\mathcal{T}_{g}$ is the Teichmüller space of marked hyperbolic or conformal structures on $S$. The mapping class group of $S$ will be denoted $\operatorname{Mod}_{g}$, and $\mathcal{M}_{g}=\mathcal{T}_{g}/\operatorname{Mod}_{g}$ is its moduli space.
2.1. The horocycle flow
Given a Riemann surface $X\in\mathcal{T}_{g}$, complex deformation theory identifies the cotangent space $T_{X}^{*}\mathcal{T}_{g}$ with the vector space of holomorphic quadratic differentials $Q(X)$ on $X$. Let $\mathcal{QT}_{g}$ denote the bundle of non-zero quadratic differentials over $\mathcal{T}_{g}$ and $\mathcal{QM}_{g}=\mathcal{QT}_{g}/\operatorname{Mod}_{g}$. By integrating ±$\pm\sqrt{q}$, each quadratic differential $q$ defines a flat metric on $q\setminus Z(q)$, where $Z(q)$ is the set of zeros of $q$. The area is given by $\int|q|$, and we denote by $\mathcal{Q}^{1}\mathcal{T}_{g}$ and $\mathcal{Q}^{1}\mathcal{M}_{g}$ the unit area loci, which can be identified with the unit cotangent bundles of $\mathcal{T}_{g}$ and $\mathcal{M}_{g}$, respectively.
Both $\mathcal{QT}_{g}$ and $\mathcal{QM}_{g}$ and their unit area loci are stratified by the number and orders of zeros of the differentials. Every partition $\underline{\kappa}=(\kappa_{1},...,\kappa_{n})$ of $44g-$ 4 and $\varepsilon\in\{\pm 1\}$ determines a (possibly empty) stratum $\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)$ of quadratic differentials that are ($\varepsilon=+1$) or are not ($\varepsilon=-1$) squares of abelian differentials. Equivalently, the completion of the flat metric defined by $q\in\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)$ has $n$ cone points, each of which has cone angle excess $\kappa_{i}\pi$. The rotational holonomy of the metric lives in $\mathbb{Z}/2\mathbb{Z}$ and is trivial if and only if $\varepsilon=+1$. We use $\mathcal{QM}_{g}(\underline{\kappa};\varepsilon)$ to denote a stratum of $\mathcal{QM}_{g}$ and add a $1$ superscript to denote unit area loci. Strata are not necessarily connected, but at least at the level of moduli space their connected components are classified [KZ03, Lan08, CM14].
Strata are complex orbifolds of dimension $(2g-2)+n+\delta$, where $\delta=1$ if $\varepsilon=+1$ and $\delta=0$ otherwise. This can be seen via local period coordinates, constructed as follows. Every $q$ has a a holonomy double cover $\widehat{q}\to q$, branched over the zeros of odd order, so that $\widehat{q}$ is a (possibly disconnected) abelian differential with a holomorphic covering involution. Let $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ denote the $(-1)$-subspace for this involution, where $(\widehat{S},\widehat{Z})\cong(\widehat{q},Z(\widehat{q}))$. Take a triangulation of $q$ by saddle connections, singularity-free geodesics joining the cone points of $q$; this can be lifted to a triangulation $\widehat{\mathsf{T}}$ of $\widehat{q}$ by saddle connections. Integrating $\widehat{q}$ over the edges of $\widehat{\mathsf{T}}$ defines a cohomology class in $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$, and any small deformation of all of the triangles subject to the constraints imposed by $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ defines a new quadratic differential in the same stratum. The period vector mapping into $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}$ near $q$ is called a period coordinate chart. Two overlapping period coordinate charts differ by an integer linear transformation. We refer the reader to [Zor06], [Wri20], and [AM24] for a more thorough discussion of period coordinates and background on flat surfaces.
In particular, for each $q\in\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)$, the tangent space may be locally modeled by
$T_{q}\mathcal{QT}_{g}(\underline{\kappa};\varepsilon)\cong H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}.$
The same is true for $\mathcal{QM}_{g}(\underline{\kappa};\varepsilon)$ modulo orbifold issues (which can be resolved by passing to a finite cover).
There is an action of $\mathsf{SL}_{2}\mathbb{R}$ on $\mathcal{Q}^{1}\mathcal{M}_{g}$ preserving each stratum component $\mathcal{Q}$. For $M\in\mathsf{SL}_{2}\mathbb{R}$ and $q\in\mathcal{Q}$, $M\cdot q\in\mathcal{Q}$ is defined by postcomposing charts for the singular flat metric obtained by integration with the $\mathbb{R}$-linear map $M:\mathbb{C}\to\mathbb{C}$. Since these charts are only defined up to ±$\pm 1$, this $\mathsf{SL}_{2}\mathbb{R}$ action factors through $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$. In a period coordinate chart $H^{1}(\widehat{S},\widehat{Z};\mathbb{C})^{-}\cong\mathbb{C}\otimes H^{1}(\widehat{S},\widehat{Z};\mathbb{R})^{-}$, the action is given by applying $M$ to $\mathbb{C}$. The one-parameter subgroups
$A=\left\{g_{t}=\begin{pmatrix}e^{t}&\\ &e^{-t}\end{pmatrix}\mid t\in\mathbb{R}\right\}\hskip 10.00002pt\text{ and }\hskip 10.00002ptU=\left\{u_{s}=\begin{pmatrix}1&s\\ &1\end{pmatrix}\mid t\in\mathbb{R}\right\}$
induce the Teichmüller geodesic flow and unstable Teichmüller horocycle flow, respectively We denote by $P\leq\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$ the group of upper-triangular matrices generated by geodesic and horocycle flows.
On each component $\mathcal{Q}$ of each stratum of quadratic differentials, there is an $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant ergodic Masur–Smillie–Veech probability measure $\nu_{\mathcal{Q}}$ in the class of Lebesgue. It is constructed from the natural affine measure in period coordinate charts by assigning to a set $E\subset\mathcal{Q}$ the affine measure of the cone $\{tE\mid t\in(0,1]\}$. A characterization of the $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant ergodic probability measures on $\mathcal{Q}^{1}\mathcal{M}_{g}$ was given by Eskin–Mirzakhani [EM18]. See §4.2 and Theorem 4.2 below.
Differentials and foliations
Every $q\in\mathcal{QM}_{g}$ is equipped with an imaginary foliation $\operatorname{Im}(q)$ by horizontal lines (which supports a transverse measure recording the total variation of the imaginary part of $q$) as well as a real foliation $\operatorname{Re}(q)$ by vertical lines supporting a transverse measure locally of the form $|dx|$. Let $\mathcal{MF}_{g}$ denote Thurston’s space of equivalence classes of measured foliations (where two foliations are equivalent if they differ by Whitehead moves and isotopy). An important result of Gardiner–Masur building on work of Hubbard–Masur shows that the map recording the equivalence classes of $\operatorname{Im}(q)$ and $\operatorname{Re}(q)$
(1) $\mathcal{QT}_{g}\to\mathcal{MF}_{g}\times\mathcal{MF}_{g}\setminus\Delta$
is a homeomorphism [GM91, HM79], where
$\Delta:=\{(\eta,\lambda)\mid\exists\gamma\in\mathcal{MF}_{g}\text{ such that }i(\gamma,\eta)+i(\gamma,\lambda)=0\}.$
Following [Vee86] and [For02] (see also [ABEM12, Theorem 3.15]), we define the strong stable and strong unstable foliations for the Teichmüller geodesic flow as follows. At every point $q$ the strong stable (unstable) leaf through $q$ locally consists of those $q^{\prime}$ with the same real (imaginary) foliation as $q$. Each unstable leaf
$\mathcal{W}^{uu}(\lambda):=\{q\in\mathcal{QT}_{g}:\text{ the imaginary foliation }\operatorname{Im}(q)\text{ is measure equivalent to }\lambda\}$
therefore has a stratified real-analytic structure coming from restricting the linear structure of strata of $\mathcal{QT}_{g}$ to intersections with $\mathcal{W}^{uu}(\lambda)$. For every $\lambda$, define $\mathcal{MF}(\lambda):=\{\eta\in\mathcal{MF}_{g}\mid(\eta,\lambda)\notin\Delta\}.$ Using (1), we have
$\mathcal{MF}(\lambda)\cong\mathcal{W}^{uu}(\lambda)\subset\mathcal{QT}_{g},$
which we can use to equip $\mathcal{MF}(\lambda)$ with a stratified real-analytic structure.
2.2. The earthquake flow
By the uniformization theorem, $\mathcal{T}_{g}$ can also be identified with the space of marked hyperbolic structures on $S$. This viewpoint yields global analytic Fenchel–Nielsen coordinates on $\mathcal{T}_{g}$ adapted to any pair of pants decomposition $\gamma$, comprised of $33g-$ 3 coordinate length functions $\ell_{i}$ and $33g-$ 3 real valued twist functions $\tau_{i}$. A choice of Fenchel–Nielsen coordinates requires a choice of what “zero twist” should mean, however, the differentials $d\tau_{i}$ are defined. In other words, the twisting data form a principal $\mathbb{R}^{3g-3}$-bundle over the length data.
Wolpert proved that the symplectic $2$-form
$\omega_{\operatorname{WP}}=\sum_{i=1}^{3g-3}d\ell_{i}\wedge d\tau_{i}$
coincides with the imaginary part of the Weil–Petersson Kähler form, hence is invariant by $\operatorname{Mod}_{g}$ [Wol85]. The top wedge power of $\omega_{\operatorname{WP}}$ defines a volume form and corresponding Weil–Petersson measure $\operatorname{vol}_{\operatorname{WP}}$ on $\mathcal{M}_{g}$. This measure is in the class of Lebesgue and has finite total mass.
For any $X\in\mathcal{T}_{g}$, Thurston introduced the space $\mathcal{ML}(X)$ of measured geodesic laminations as a natural completion of the set of simple closed geodesics on $X$ [Thu88, CB88, PH92]. A geodesic lamination $\lambda\subset X$ is a closed set that is foliated by disjoint, simple, complete geodesics. The metric completion $X\setminus\lambda$ of its complement is a finite area hyperbolic surface with (possibly non-compact) totally geodesic boundary. A transverse measure on $\lambda$ is the assignment of a measure $\mu_{k}$ to each $C^{1}$ arc $k$ transverse to $\lambda$, such that the support of $\mu_{k}$ is $k\cap\lambda$, and such that this assignment is natural under inclusion and invariant under transverse isotopy. Throughout, we will write $\lambda$ to refer both to a transverse measure and its supporting geodesic lamination.
Neither the set of geodesic laminations nor the space $\mathcal{ML}(X)$ depends on the metric $X$; as such, we use the notation $\mathcal{ML}_{g}$ to refer to the space of measured laminations. The theory of train tracks gives a family of coordinate charts specifying polyhedral cones in $\mathcal{ML}_{g}$; in each chart, $\mathbb{N}$-weighted multicurves specify an integer lattice. Transition maps between these charts are in $\mathsf{SL}_{n}$, and hence this lattice can be used to define a $\operatorname{Mod}_{g}$-invariant Lebesgue-class measure on $\mathcal{ML}_{g}$ called the Thurston measure $\mu_{\operatorname{Th}}$. The generic $\lambda$ with respect to $\mu_{\operatorname{Th}}$ is maximal, in that $X\setminus\lambda$ is a union of ideal triangles.
Given $X\in\mathcal{T}_{g}$, the function that sends a multicurve $\gamma$ to its geodesic length on $X$ extends homogeneously and continuously to the space of measured laminations. Denote by $\mathcal{ML}^{1}(X)$ the sphere of unit length measured laminations on $X$; then $\mu_{\operatorname{Th}}$ defines a finite measure $\hat{\mu}_{\operatorname{Th}}^{X}$ on $\mathcal{ML}^{1}(X)$ by
$\hat{\mu}_{\operatorname{Th}}^{X}(E):=\mu_{\operatorname{Th}}\{c\lambda\mid\lambda\in E,c\in[0,1]\}.$
Let $B(X)$ denote the mass of $\hat{\mu}_{\operatorname{Th}}^{X}$, equivalently, the $\mu_{\operatorname{Th}}$ measure of the ball of length $\leq 1$ laminations on $X$.
There is a flat bundle $\mathcal{P}^{1}\mathcal{T}_{g}$ over $\mathcal{T}_{g}$ with fiber $\mathcal{ML}^{1}(X)$ over $X$; the quotient by $\operatorname{Mod}_{g}$ is denoted by $\mathcal{P}^{1}\mathcal{M}_{g}$. By [Thu98, Theorem 5.1], these bundles can also be identified with the unit cotangent bundles of $\mathcal{T}_{g}$ and $\mathcal{M}_{g}$, respectively. Integrating the fiberwise measures $\hat{\mu}_{\operatorname{Th}}^{X}$ against the Weil–Petersson volume yields the $\operatorname{Mod}_{g}$-invariant, Lebesgue-class Mirzakhani measure on $\mathcal{P}^{1}\mathcal{T}_{g}$. Its local pushforward to $\mathcal{P}^{1}\mathcal{M}_{g}$ is a finite measure $\mu_{\operatorname{Mirz}}$ that we also call by the same name. The total total mass of $\mu_{\operatorname{Mirz}}$ is
$b_{g}:=\int_{\mathcal{M}_{g}}B(X)\,d\operatorname{vol}_{\operatorname{WP}}(X),$
and by construction, its pushforward to $\mathcal{M}_{g}$ under the forgetful map is $B(X)d\operatorname{vol}_{\operatorname{WP}}$(X).
The Mirzakhani measure is invariant and ergodic for the (right) earthquake flow, defined as follows. Given a $(X,c\gamma)\in\mathcal{P}^{1}\mathcal{M}_{g}$, where $c\gamma$ is a weighted simple closed curve, $\operatorname{Eq}_{s}(X,c\gamma)=(\operatorname{Eq}_{s}(X,c\gamma),c\gamma)$, where the first coordinate is the hyperbolic metric obtained by cutting $X$ open along $\gamma$ and re-gluing with a right twist of length $cs$. Equivalently, this flow translates by $cs$ in Fenchel–Nielsen coordinates adapted to a pants decomposition containing $\gamma$. For $\lambda\in\mathcal{ML}(S)$ approximated by weighted simple closed curves $c_{n}\gamma_{n}$, the homeomorphisms $X\mapsto\operatorname{Eq}_{s}(X,c_{n}\gamma_{n})$ converge to a map called the right earthquake in $\lambda$ (see [Ker83] or [Thu06]). This map also admits an independent description by cutting and re-gluing all of the complementary components of $\lambda$ on $X$ with a rightwards shear, moving all components relative to one another by amounts dictated by the transverse measure to $\lambda$.
2.3. The conjugacy
There is an important connection between the horocycle and earthquake flows introduced above: they are Borel isomorphic (see Theorem 2.3 below). This relationship was first observed by Mirzakhani on the full $\mu_{\operatorname{Mirz}}$-measure subset of pairs $(X,\lambda)\in\mathcal{P}^{1}\mathcal{M}_{g}$ where $\lambda$ is a maximal geodesic lamination [Mir08] and was extended to the entirety of $\mathcal{P}^{1}\mathcal{M}_{g}$ by the authors [CF24b].
This isomorphism is constructed by assigning a “dual” foliation to every lamination $\lambda$ on a hyperbolic surface $X$. Consider a component $Y$ of $X\setminus\lambda$. Most points of $Y$ have a unique closest point on $\partial Y$; those that have multiple closest points on $\partial Y$ form a piecewise geodesic $1$-complex $\operatorname{\mathsf{Sp}}(Y)$ called the spine of $Y$ (see also §3.2 below). The fibers of the closest point–projection map $Y\setminus\operatorname{\mathsf{Sp}}(Y)\to\partial Y$ form a foliation of $Y\setminus\operatorname{\mathsf{Sp}}(Y)$. Extending continuously across $\operatorname{\mathsf{Sp}}(Y)$ yields a piecewise geodesic singular foliation $\mathcal{O}_{\partial Y}(Y)$ called the orthogeodesic foliation. This foliation has $k$-pronged singularities at points that are equidistant from $k$ different points of $\partial Y$, and every endpoint of every leaf of $\mathcal{O}_{\partial Y}(Y)$ meets $\partial Y$ orthogonally.
Because their leaves are always orthogonal to $\lambda$, the orthogeodesic foliations on the components of $X\setminus\lambda$ can be extended continuously across $\lambda$, yielding a singular foliation $\mathcal{O}_{\lambda}(X)$ on $S$. This foliation also carries a natural transverse measure, defined by assigning to any geodesic arc in $\lambda$ its Lebesgue measure and then extending uniquely to all arcs via transverse isotopy. See [CF24b, Section 5] for more details.
This defines a map $\mathcal{O}_{\lambda}:\mathcal{T}_{g}\to\mathcal{MF}_{g}$ recording the equivalence class of $\mathcal{O}_{\lambda}(X)$; the following theorem records the global properties of $\operatorname{\mathcal{O}_{\lambda}}$. When $\lambda$ is maximal, it is due to Thurston [Thu98]; see also [Bon96] for a treatment in terms of transverse Hölder distributions. For arbitrary $\lambda$, the following is a direct consequence of Theorems D and E of [CF24b].
Theorem 2.1.
For every $\lambda\in\mathcal{ML}_{g}$, the map $\operatorname{\mathcal{O}_{\lambda}}$ is a stratified real-analytic homeomorphism
$\operatorname{\mathcal{O}_{\lambda}}:\mathcal{T}_{g}\to\mathcal{MF}(\lambda)\cong\mathcal{W}^{uu}(\lambda)$
that is equivariant with respect to the stabilizer of $\lambda$ in $\operatorname{Mod}_{g}$ and satisfies $i(\mathcal{O}_{\lambda}(X),\lambda)=\ell_{X}(\lambda)$.
Remark 2.2.
Theorem D of [CF24b] establishes that $\operatorname{\mathcal{O}_{\lambda}}$ is a homeomorphism. Theorem E therein asserts that $\operatorname{\mathcal{O}_{\lambda}}$ factors as $\operatorname{\mathcal{O}_{\lambda}}=I_{\lambda}^{-1}\circ\sigma_{\lambda}$, where $\sigma_{\lambda}$ is stratified real-analytic and $I_{\lambda}^{-1}$ is linear when restricted to each stratum. Thus $\operatorname{\mathcal{O}_{\lambda}}$ is also stratified real-analytic.
It is well known that $\mathcal{ML}_{g}$ and $\mathcal{MF}_{g}$ are canonically homeomorphic [Lev83]. Combining this fact with Theorem 2.1 and (1), we obtain a $\operatorname{Mod}_{g}$-equivariant bijection
$\begin{array}[]{rcccl}\mathcal{O}:\mathcal{PT}_{g}=&\mathcal{T}_{g}\times\mathcal{ML}_{g}&\to&\mathcal{MF}_{g}\times\mathcal{MF}_{g}\setminus\Delta&\cong\mathcal{QT}_{g}.\\ &(X,\lambda)&\mapsto&(\operatorname{\mathcal{O}_{\lambda}}(X),\lambda)\end{array}$
Moreover, for $(X,\lambda)\in\mathcal{PT}_{g}$, the real foliation of $\mathcal{O}(X,\lambda)$ is actually isotopic to $\mathcal{O}_{\lambda}(X)$ (see [CF24b, §5.3]).
The fact that the length of $\lambda$ on $X$ is equal to the area of $\mathcal{O}(X,\lambda)$ combined with the $\operatorname{Mod}_{g}$-equivariance of $\mathcal{O}$ imply that it induces a bijection
$\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$
of the same name. The following omnibus theorem combines the main theorems of [Mir08] and [CF24b] together with a special case of the main result of [CF24a].
Theorem 2.3.
The map $\mathcal{O}:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ is a bijective Borel isomorphism [^3] Theorem C of [CF24b] and Theorem 15.4 of [CF24a]. with the following properties:
  1. (1)
    $\mathcal{O}$ takes earthquake flow to horocycle flow in a time-preserving way: $\mathcal{O}\circ\operatorname{Eq}_{s}=u_{s}\circ\mathcal{O}$ [^4] Theorem A of [CF24b]; see also Theorem 1.1 of [Mir08].
  2. (2)
    $\mathcal{O}_{*}\mu_{\operatorname{Mirz}}/b_{g}$ is the Masur–Smillie–Veech probability measure on the principal stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$. [^5] Section 1.2 of [Mir08].
  3. (3)
    Each $k$-pronged singularity of $\operatorname{\mathcal{O}_{\lambda}}(X)$ corresponds to an order $(k-2)$ zero of $\mathcal{O}(X,\lambda)$ and vice versa. [^6] Proposition 5.10 of [CF24b].
  4. (4)
    If $\mathcal{Q}$ is a component of a stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$ and $q\in\mathcal{Q}$ has no horizontal saddle connections, then $q$ is a point of continuity for $\mathcal{O}^{-1}|_{\mathcal{Q}}$. [^7] A stronger version of this is proved as Theorem 15.6 of [CF24a].
All $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant probability measures on strata give zero measure to the set of differentials with horizontal saddle connections (see §4.2 below). Item (4) therefore allows us to pull back equidistribution to such measures [Bil68, Theorem 5.1], and implies Theorem 1.4 from the Introduction.
Remark 2.4.
Mirzakhani pointed out that $\mathcal{O}$ (or any extension of her original conjugacy) is discontinuous along sequences $(X_{n},\lambda_{n})$ where the measured and Hausdorff limits of maximal laminations $\lambda_{n}$ disagree [Mir08, p. 33]. It is now known that there can be no continuous map $\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{Q}^{1}\mathcal{M}_{g}$ conjugating the earthquake and horocycle flows [AHW23], so Theorem 2.3 (and the more general Theorem A of [CF24a]) are essentially optimal.
The conjugacy $\mathcal{O}$ allows us to pull back the Teichmüller geodesic flow from $\mathcal{Q}^{1}\mathcal{M}_{g}$ to define a measurable dilation flow $(\operatorname{Dil}_{t})$ on $\mathcal{P}^{1}\mathcal{M}_{g}$ that normalizes the earthquake flow (see [CF24b, §15.3]):
$\operatorname{Dil}_{t}\circ\operatorname{Eq}_{s}\circ\operatorname{Dil}_{-t}=\operatorname{Eq}_{e^{2t}s}.$
This transformation can also be characterized by scaling the orthogeodesic foliation (and hence the spine of $X\setminus\lambda$). Given any $X\in\mathcal{T}_{g}$ and $\lambda\in\mathcal{ML}_{g}$, the dilation ray $\{X_{t}\}$ directed by $\lambda$ is the projection of $\operatorname{Dil}_{t}(X,\lambda)$ to $\mathcal{T}_{g}$. By Theorem 2.1, it is uniquely specified by the equation $\operatorname{\mathcal{O}_{\lambda}}(X_{t})=e^{t}\operatorname{\mathcal{O}_{\lambda}}(X).$
3. Hyperbolic twist tori
We begin by establishing some notation for twist tori and associated measures on moduli spaces. Throughout this section, we use Conjecture 1.1 as a motivating example but phrase all of our results in generality.
3.1. Twist tori
Let $\gamma=(\gamma_{1},\ldots,\gamma_{n})$ be a multicurve on $S$; throughout the paper we use $S\setminus\gamma$ to denote the (possibly disconnected) closed surface obtained by removing an open annulus about each $\gamma_{i}$. Associated to $\gamma$ there is a cut-and-glue fibration that expresses a hyperbolic structure on $S$ in terms of the induced hyperbolic structure on $S\setminus\gamma$ and how these structures are glued together. Let $\mathcal{T}(S\setminus\gamma)$ denote the space of marked [^8] Here, the marking does not record twisting about the boundary. hyperbolic structures with totally geodesic boundary on $S\setminus\gamma$ so that for each $i$, the two boundaries of $S\setminus\gamma$ that are glued along $\gamma_{i}$ have the same length. Then we have the following principal $\mathbb{R}^{n}$-bundle:
${\mathbb{R}^{n}}$${\mathcal{T}_{g}}$${\mathcal{T}(S\setminus\gamma)}$$\scriptstyle{\text{cut}_{\gamma}}$
where the $\mathbb{R}^{n}$ fibers record the twisting about the curves of $\gamma$. In the case that $\gamma$ is a pair of pants, these are just the usual Fenchel–Nielsen coordinates. For an arbitrary measured lamination $\lambda$, the authors developed shear-shape coordinates that parametrize $\mathcal{T}_{g}$ in a similar way [CF24b].
[…]
$\text{cut}_{\gamma}^{-1}(Y)/\langle T_{\gamma_{i}}\rangle\cong(\mathbb{R}/\ell_{1}\mathbb{Z})\times\ldots\times(\mathbb{R}/\ell_{n}\mathbb{Z})\subset\mathcal{T}_{g}/\langle T_{\gamma_{i}}\rangle$
equipped with a measure $\widetilde{\tau}_{\gamma}(Y)$ of mass $\prod\ell_{i}$(which is the local pushforward of Lebesgue).
Let $\pi:\mathcal{T}_{g}\to\mathcal{M}_{g}$ and $\tilde{\pi}:\mathcal{T}_{g}/\langle T_{\gamma_{i}}\rangle\to\mathcal{M}_{g}$ denote the natural (orbifold) covering maps.
Definition 3.1.
The twist torus associated to $Y\in\mathcal{T}(S\setminus\gamma)$ is
$\mathbb{T}_{\gamma}(Y):=\pi\left(\text{cut}_{\gamma}^{-1}(Y)\right)=\tilde{\pi}\left(\text{cut}_{\gamma}^{-1}(Y)/\langle T_{\gamma_{i}}\rangle\right)\subset\mathcal{M}_{g}.$
The associated twist torus measure on $\mathcal{M}_{g}$ is $\tau_{\gamma}(Y):=\tilde{\pi}_{*}\left(\widetilde{\tau}_{\gamma}(Y)\right).$
Lifting to the cotangent bundle
In order to apply the results of [CF24a] and connect hyperbolic twist tori to their flat counterparts, we will need to first lift twist tori to $\mathcal{P}^{1}\mathcal{M}_{g}$. To do so, choose any weight vector $\mathbf{h}=(h_{1},\ldots,h_{n})\in\mathbb{R}_{>0}^{n}$ satisfying $\bm{\ell}\cdot\mathbf{h}=1$. We can then lift the fiber $\text{cut}_{\gamma}^{-1}(Y)$ into $\mathcal{P}^{1}\mathcal{T}_{g}$ by setting the second coordinate equal to ${\mathbf{h}}\gamma$, the weighted multicurve ${h_{1}}\gamma_{1}\cup\ldots\cup{h_{n}}\gamma_{n}$. Quotienting out by the twist subgroup yields an embedded torus in $\mathcal{P}^{1}\mathcal{T}_{g}/\langle T_{\gamma_{i}}\rangle$ with its Lebesgue measure, and pushing down to moduli space we get a torus
$\mathcal{P}\mathbb{T}_{\gamma}(Y,\mathbf{h}):=\pi\left(\text{cut}_{\gamma}^{-1}(Y)\times\{\mathbf{h}\gamma\}\right)\subset\mathcal{P}^{1}\mathcal{M}_{g}$
where $\pi$ is again the quotient by the action of the mapping class group.
We can now define a measure supported on $\mathcal{P}\mathbb{T}_{\gamma}(Y,\mathbf{h})$ similarly to how we obtained $\tau_{\gamma}(Y)$: there is a natural Lebesgue measure on $\text{cut}_{\gamma}^{-1}(Y)\times\{\mathbf{h}\gamma\}$, whose local pushforward yields a finite measure after quotienting by the twist subgroup, and we define $\mu_{\gamma}(Y,\mathbf{h})$ to be the further pushforward of this measure to $\mathcal{P}^{1}\mathcal{M}_{g}$, normalized to have unit mass. Unpacking definitions, we see that
(2) $p_{*}\left(\mu_{\gamma}(Y,\mathbf{h})\right)=\frac{\tau_{\gamma}(Y)}{\ell_{1}\cdots\ell_{n}}$
as measures on $\mathcal{M}_{g}$, where $p:\mathcal{P}^{1}\mathcal{M}_{g}\to\mathcal{M}_{g}$ is the natural projection map.
Remark 3.2.
The map $p$ does not necessarily restrict to a homeomorphism between $\mathbb{T}_{\gamma}(Y)$ and $\mathcal{P}\mathbb{T}_{\gamma}(Y,\mathbf{h})$; it is generally just an (orbifold) covering. For example, if $\gamma=(\gamma_{1},\gamma_{2})$ where each $\gamma_{i}$ is nonseparating but $h_{1}\neq h_{2}$, then the stabilizer of $\gamma$ is larger than the stabilizer of $\mathbf{h}\gamma$.
3.2. Arc systems and ribbon graphs
In order to discuss families of expanding twist tori, we will need to vary the hyperbolic structure $Y$ on the complement of $\gamma$. This is easy when $\gamma$ is a pants decomposition, because one can just scale up the lengths of the pants curves. In general, we will consider the pushforwards of $\mathcal{P}\mathbb{T}_{\gamma}(Y,\mathbf{h})$ by the dilation flow (defined in §2.3 above).
To describe these tori more concretely, we recall from Section 2.3 that the spine $\operatorname{\mathsf{Sp}}(Y)$ of a hyperbolic surface with boundary $Y$ is the set of points that have multiple closest points on $\partial Y$. The spine is naturally a ribbon graph (i.e., a graph with cyclic orderings at its vertices) whose vertices of valence $k$ correspond to $k$-pronged singularities of $\mathcal{O}_{\partial Y}(Y)$. It is also equipped with a metric given by assigning to each edge of $\operatorname{\mathsf{Sp}}(Y)$ the length of either of its closest point–projections to $\partial Y$; equivalently, this is the mass of the transverse measure deposited by the orthogeodesic foliation on the geodesic connecting the two vertices.
Dually, the data of $\operatorname{\mathsf{Sp}}(Y)$ can also be recorded as a weighted, filling arc system on $Y$. Leaves of the orthogeodesic foliation $\mathcal{O}_{\partial Y}(Y)$ break up into finitely many isotopy classes of arcs rel $\partial Y$ and we use $\underline{\alpha}$ to denote the set of these isotopy classes. Assigning to each arc the length of its dual edge of $\operatorname{\mathsf{Sp}}(Y)$ then defines a positive weight system on $\underline{\alpha}$. See [CF24b, §6] for a more thorough discussion of this duality.
The map taking a taking a hyperbolic surface with boundary to its spine (or dual arc system) in fact defines a stratified piecewise real-analytic $\operatorname{Mod}_{g,b}$–equivariant homeomorphism [Luo07]
$\operatorname{\mathsf{Sp}}:\mathcal{T}_{g,b}\xrightarrow{\sim}\mathcal{TRG}_{g,b},$
where $\mathcal{T}_{g,b}$ is the space of marked hyperbolic surfaces with genus $g$ and $b$ boundaries and $\mathcal{TRG}_{g,b}$ is the space of marked metric ribbon graphs (up to homotopy) with the same genus and number of boundaries. [^9] A marking of a ribbon graph $\Gamma$ is a choice of deformation retract $S_{g,b}\to\Gamma$. See also [Ush99, Do08, Mon09].
The spine map takes surfaces with boundary lengths $\bm{\ell}$ to ribbon graphs whose boundary lengths are also $\bm{\ell}$. Given a multicurve $\gamma$ on a closed surface $S$, we can therefore piece together the spine maps on each component of $S\setminus\gamma$ to get a homeomorphism
(3) $\operatorname{\mathsf{Sp}}:\mathcal{T}(S\setminus\gamma)\xrightarrow{\sim}\mathcal{TRG}(S\setminus\gamma)$
where $\mathcal{TRG}(S\setminus\gamma)$ is defined analogously to $\mathcal{T}(S\setminus\gamma)$; compare [AHC22, pg. 39]. We observe that when $\gamma$ is a pants decomposition, $\operatorname{\mathsf{Sp}}(Y)$ is completely determined by the boundary lengths $\bm{\ell}$. In the general case, the lengths of the $\gamma_{i}$ are explicitly obtainable from the metric on $\operatorname{\mathsf{Sp}}(Y)$, and we will continue to use $\bm{\ell}\in\mathbb{R}^{n}$ to denote the vector of lengths of the curves of $\gamma$.
Notation.
In light of (3), throughout the rest of the paper we will refer to twist tori and their measures in terms of the complementary metric ribbon graph $\mathsf{Y}=\operatorname{\mathsf{Sp}}(Y)\in\mathcal{TRG}(S\setminus\gamma)$. For example,
$\mu_{\gamma}(\mathsf{Y}):=\mu_{\gamma}(Y)\text{ and }\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h}):=\mathcal{P}\mathbb{T}_{\gamma}(Y,\mathbf{h}).$
The discussion in this section should make it clear how a twist torus changes when it is pushed forward by the dilation flow: the spine of the complementary subsurface is rescaled. That is,
(4) $\operatorname{Dil}_{t}\left(\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})\right)=\mathcal{P}\mathbb{T}_{\gamma}(e^{t}\mathsf{Y},e^{-t}\mathbf{h}).$
Walls and chambers
The space $\mathcal{TRG}(S\setminus\gamma)$ naturally has a piecewise polyhedral structure, in which top-dimensional faces correspond to trivalent metric ribbon graphs and lower-dimensional ones correspond to ribbon graphs with higher valence vertices (this combinatorial structure is also readily apparent in the dual picture in terms of weighted filling arc systems). For a given ribbon graph $\Gamma$ of the same topological type as $S\setminus\gamma$, we set $F(\Gamma)$ to denote the facet of $\mathcal{TRG}(S\setminus\gamma)$ consisting of metrics on $\Gamma$ satisfying the gluing conditions. Luo’s work [Luo07] implies that $\operatorname{\mathsf{Sp}}$ restricts to a real-analytic homeomorphism between each facet and a corresponding subvariety of $\mathcal{T}(S\setminus\gamma)$.
It is important to note that that the dimension of $F(\Gamma)$ is not necessarily the number of edges of $\Gamma$. For the same reason, some choices of $\Gamma$ do not correspond to (nonempty) facets: this is because combinatorial conditions on the spine can enforce constraints on the lengths of the boundaries, which might not be compatible with the gluing conditions over the curves of $\gamma$. See Figure 3 for an example of this phenomenon.
Figure 3. A pair of spines that does not appear in $\mathcal{TRG}(S\setminus\gamma)$ and their dual arc systems. Together, the combinatorial conditions imply that the length of $c$ must be 0, which cannot happen inside $\mathcal{T}(S\setminus\gamma)$.
3.3. Horospheres and slices
Given any $\mathbf{h}=(h_{1},\ldots,h_{n})\in\mathbb{R}_{>0}^{n}$, define the length $1$ horosphere in $\mathcal{T}_{g}$ based at the weighted multicurve ${\mathbf{h}}\gamma={h_{1}}\gamma_{1}\cup\ldots\cup{h_{n}}\gamma_{n}$ to be the $1$-level set of the (total) length function. [^10] This discussion still holds for the level sets of other functions of the length of the components of $\gamma$; compare [AH21, Liu20]. That is,
$H_{\gamma}(\mathbf{h}):=\left\{X\in\mathcal{T}_{g}\mid\sum_{i=1}^{n}h_{i}\ell_{X}(\gamma_{i})=1\right\}.$
Each horosphere has a natural lift $\mathcal{P}H_{\gamma}(\mathbf{h})$ to $\mathcal{P}^{1}\mathcal{T}_{g}$ with $\mathbf{h}\gamma$ as the second coordinate. Both the horosphere and its lift are homeomorphic to $\mathbb{R}^{6g-7}$ and are fibered by twist tori.
Following Forni [For21], we define a horospherical measure to be any probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ that is absolutely continuous with respect to the Lebesgue class with continuous density and such that its restriction to each earthquake flow line is the restriction of Lebesgue to an interval. We also say that the pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$ is horospherical.111111 For example, in [AH21, Liu20, AHC22], such measures are obtained by taking certain $\operatorname{Stab}(\mathbf{h}\gamma)$-invariant measures on $\mathcal{P}H_{\gamma}(\mathbf{h})$, taking a local pushforward to the intermediate quotient $\mathcal{P}^{1}\mathcal{T}_{g}/\operatorname{Stab}(\mathbf{h}\gamma)$, then pushing down to $\mathcal{P}^{1}\mathcal{M}_{g}$.
Following our discussion of the wall-and-chamber structure of $\mathcal{TRG}(S\setminus\gamma)$, given any facet $F(\Gamma)$, we also define the corresponding length $1$ horosphere slice as follows:
$H_{\gamma}(\mathbf{h},\Gamma):=\left\{X\in\mathcal{T}_{g}\mid\sum_{i=1}^{n}h_{i}\ell_{X}(\gamma_{i})=1\text{ and }\operatorname{\mathsf{Sp}}(X\setminus\gamma)\in F(\Gamma)\right\}.$
As before, there is a natural lift to $\mathcal{P}^{1}\mathcal{T}_{g}$, both the horosphere slice and its lift are homeomorphic to a Euclidean space, and both are fibered by twist tori. We say that a probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ is a horosphere slice measure if it is in the class of Lebesgue with continuous density (now on the slice) and the conditional measures on earthquake flow lines are Lebesgue on intervals. Just as above, we also use this term to denote pushforward of such a measure to $\mathcal{P}^{1}\mathcal{M}_{g}$.
Remark 3.3.
There are natural examples of horospherical and horosphere slice measures coming from restricting the Weil-Petersson form or by integrating the natural affine measure on facets of $\mathcal{TRG}(S\setminus\gamma)$ against twist parameters. For applications, one usually wishes to fix such a choice and compute its mass (pre-normalization). For this reason, the definition of horospherical measures is usually a little more involved in order to record factors coming from orbifold issues. We refer the interested reader to [AHC22, §7] or [AH20].
4. Flat twist tori
In this section, we discuss the flat counterparts of twist tori, horospheres, slices, and the relevant measures. We prove using the strong rigidity phenomena of [EM18, EMM15] for the $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-action on strata of quadratic differentials that (the union of) such expanding families equidistribute to the affine measure on an affine invariant subvariety.
4.1. Twist tori and horospheres
Corresponding to $\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is a flat twist torus parameterizing quadratic differentials built from horizontal cylinders with core curves $\gamma$, heights $\mathbf{h}$, and lengths (circumferences) $\bm{\ell}$; these cylinders are glued with arbitrary twists along the metric ribbon graph $\mathsf{Y}$. Indeed, define
$\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})=\mathcal{O}(\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h}))\subset\mathcal{Q}^{1}\mathcal{M}_{g}$
and let
$\nu_{\gamma}(\mathsf{Y},\mathbf{h})=\mathcal{O}_{*}\mu_{\gamma}(\mathsf{Y},\mathbf{h})$
be the associated probability measure. By Theorem 2.1, $\mathcal{O}$ restricts to a real analytic homeomorphism between $\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ and $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$.
The differentials $q\in\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ are completely horizontally periodic, because their horizontal foliations are measure equivalent to $\mathbf{h}\gamma$. With definitions as in [Wri15] or §5.1 below, the flat twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is just the orbit of any such by the abelian group $\mathbb{R}^{\gamma}$ of twists in the horizontal cylinders, and $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$ is Lebesgue on $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. We observe that $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is locally an affine submanifold in period coordinates, cut out by equations of the form $\operatorname{Im}(z)=h_{i}$ and $z=\ell(e)$, where $e$ is an edge of $\mathsf{Y}$ with length $\ell(e)$. By our convention that $\bm{\ell}\cdot\mathbf{h}=1$, the intersection of these affine subspaces lies in the unit area locus.
Similarly, for a given horosphere $\mathcal{P}H_{\gamma}(\mathbf{h})$, we define the extremal length horosphere
$\mathcal{Q}H_{\gamma}(\mathbf{h})=\mathcal{O}(\mathcal{P}H_{\gamma}(\mathbf{h}))\subset\mathcal{Q}^{1}\mathcal{T}_{g},$
which is actually a lift of the $1$-level set of the extremal length function $L_{\mathbf{h}\gamma}:\mathcal{T}_{g}\to\mathbb{R}_{>0}$ [CF24b, GM91]. Using Theorem 2.1, $\mathcal{Q}H_{\gamma}(\mathbf{h})$ is a leaf of the strong unstable foliation. For any facet of $\mathcal{TRG}(S\setminus\gamma)$ we can also define the extremal length horosphere slice
$\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)=\mathcal{O}(\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)).$
By construction of $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$ and Theorems 2.1 and 2.3, $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ is actually a leaf of the strong unstable foliation in the ambient stratum containing it. The angle around each singularity of the flat metric is $\pi$ times the valence of the corresponding vertex of $\Gamma$; see §2 of [CF24b].
We observe how twist tori and their associated measures interact with the $P$ action.
Lemma 4.1.
For any $\gamma$, $\mathsf{Y}$, and $\mathbf{h}$, the twist torus measure $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$ is $U$-invariant and satisfies
(5) $(g_{t})_{*}\nu_{\gamma}(\mathsf{Y},\mathbf{h})=\nu_{\gamma}(e^{t}\mathsf{Y},e^{-t}\mathbf{h}).$
Proof.
The action of the horocycle flow can be written as the sum of cylinder twists on $\gamma_{i}$ of total length $h_{i}$; since Lebesgue measure on a torus is translation invariant, we see that twist torus measures are $U$-invariant.
As the Teichmüller geodesic flow stretches all horizontal data and shrinking all vertical data, it takes twist tori to twist tori. To see that the measures are taken to one another, we observe that the twist torus measure is defined via Lebesgue measure on the twists, which constitute horizontal data. Therefore, the Teichmüller geodesic flow uniformly expands the (non-normalized) Lebesgue measure on the twist torus, hence preserves the normalized Lebesgue measure. ∎
4.2. Affine invariant subvarieties
Lemma 4.1 tells us that we may view expanding twist tori as pushforwards of a single torus along the geodesic flow. By averaging these measures, we can invoke the seminal work of Eskin, Mizkakhani, and Mohammadi on the $P$- and $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant measures and orbit closures in strata of holomorphic differentials. We briefly summarize their results below.
The unit area locus $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$ of a $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant immersed sub-orbifold $\mathbb{R}\mathcal{L}\subset\mathcal{QM}_{g}$ that is locally cut out by $\mathbb{C}$-linear equations defined over $\mathbb{R}$ in period coordinates is called an affine invariant subvariety or AIS. An AIS $\mathcal{L}$ comes equipped with an $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant probability affine measure $\nu_{\mathcal{L}}$, where if $m$ is Lebesgue measure on $\mathbb{R}$, then $dm\,d\nu_{\mathcal{L}}$ is locally an affine linear measure on a finite union of complex subspaces in period coordinates; see [EM18, Definition 1.1]. While we will not use it directly, we mention that Filip proved that every AIS $\mathcal{L}$ is an algebraic variety [Fil16].
The following combines [EM18, Proposition 1.3 and Theorem 1.4] and [EMM15, Theorems 2.1 and 2.10].
Theorem 4.2.
Let $\mathcal{Q}$ be any component of any stratum of $\mathcal{Q}^{1}\mathcal{M}_{g}$.
  1. (1)
    (Measure classification): Every $P$-invariant ergodic measure on $\mathcal{Q}$ is also $\operatorname{\mathsf{PSL}}_{2}\mathbb{R}$-invariant and is the affine measure on an AIS, of which there are at most countably many.
  2. (2)
    (Orbit closures): For any $q\in\mathcal{Q}$, the orbit closure $\overline{P\cdot q}=\overline{\operatorname{\mathsf{PSL}}_{2}\mathbb{R}\cdot q}$ is an AIS.
  3. (3)
    (Genericity): For any $q\in\mathcal{Q}$, the orbit $P\cdot q$ equidistributes in its closure $\mathcal{L}$. That is, for any $f\in C_{c}(\mathcal{Q})$ and any $r\in\mathbb{R}\setminus\{0\}$,
[…]
Applying Theorem 4.2 to the setting of Lemma 4.1, we deduce that the time-averaged pushforwards of a twist torus equidistribute to its $P$–orbit closure.
Proposition 4.3.
For any multicurve $\gamma$ on $S$, any ribbon graph spine $\mathsf{Y}\in\mathcal{TRG}(S\setminus\gamma)$, and any cylinder heights $\mathbf{h}\in\mathbb{R}_{>0}^{n}$ with $\bm{\ell}\cdot\mathbf{h}=1$, we have that
[…]
as $T\to\infty$, where $\nu_{\mathcal{L}}$ is the unique affine probability measure supported on $\mathcal{L}:=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})}.$
Proof.
Since $\gamma$, $\mathsf{Y}$, and $\mathbf{h}$ are fixed, throughout the proof we will denote $g_{t}\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ by $\mathcal{Q}\mathbb{T}^{t}$ and the associated twist torus measure on $\mathcal{Q}\mathbb{T}^{t}$ by $\nu^{t}$.
We first consider the (generic) situation in which the moduli $h_{i}/\ell_{i}$ of the horizontal cylinders are linearly independent over $\mathbb{Q}$. In this case, the torus $\mathcal{Q}\mathbb{T}^{0}$ is a minimal set for the horocycle flow and so the $U$ orbit closure of any point is the entire torus. We therefore see that
$\overline{P\cdot q}=\overline{P\cdot\mathcal{Q}\mathbb{T}^{0}}=\mathcal{L}$
for every $q\in\mathcal{Q}\mathbb{T}^{0}.$
Invoking Theorem 4.2, this implies that the Fölner sets $\{g_{t}u_{s}q:t\in[0,T],s\in[0,r]\}$ based at any $q\in\mathcal{Q}\mathbb{T}^{0}$ equidistribute to $\nu_{\mathcal{L}}$ as $T\to\infty$. Integrating against the twist torus measure and applying Lemma 4.1 yields
[…]
for every $f\in C_{c}(\mathcal{Q})$, where $\mathcal{Q}$ is the ambient stratum. This is what we wanted to show.
In general, the twist torus $\mathcal{Q}\mathbb{T}^{0}$ may not be minimal for the horocycle flow. In this case, we show that the $P$–orbit closure of $\nu^{0}$–almost every point is $\mathcal{L}$; the rest of the argument then proceeds as above.
On any $q\in\mathcal{Q}\mathbb{T}^{0}$, choose a basis for relative homology consisting of horizontal saddle connections and saddles which connect the boundary components of the cylinders of $q$; this gives a local period coordinate chart for the stratum $\mathcal{Q}$. Fix a finite collection121212 In fact, one can cover a lift of the twist torus to $\mathcal{Q}^{1}\mathcal{T}_{g}$ with a single chart of this form, but we use a finite collection of restricted charts so that we do not need to worry about the $\operatorname{Stab}(\gamma)$ action. of such charts that covers $\mathcal{Q}\mathbb{T}^{0}$. Let $\mathcal{N}\subsetneq\mathcal{L}$ be any affine invariant subvariety not containing $\mathcal{Q}\mathbb{T}^{0}$. In any of the period charts constructed above, $\mathcal{Q}\mathbb{T}^{0}$ is cut out by affine equations of the form $\operatorname{hol}(s)=\ell_{i}\in\mathbb{R}$ (specifying the length of the horizontal saddle connections) and equations of the form $\operatorname{Im}(\operatorname{hol}(s))=h_{i}$ (specifying the heights of the cylinders); compare §4.1. On the other hand, in period coordinates $\mathcal{N}$ is a union of finitely many linear subspaces, each of which is cut out by linear equations with real coefficients. Therefore, we see that in each chart, $\mathcal{Q}\mathbb{T}^{0}\cap\mathcal{N}$ is a union of finitely many proper subspaces, hence has $\nu^{0}$ measure 00. Moreover, we note that each subspace must be tangent to a proper subtorus of $\mathcal{Q}\mathbb{T}^{0}$, else its closure would be the entirety of $\mathcal{Q}\mathbb{T}^{0}$.
As the $P$–orbit closure of any $q$ is always an AIS and there are only countably many such (Theorem 4.2.1), we see that
$\{q\in\mathcal{Q}\mathbb{T}^{0}:\overline{Pq}\neq\mathcal{L}\}$
is a union of countably many proper subtori of $\mathcal{Q}\mathbb{T}^{0}$. In particular, this set has measure zero with respect to $\nu^{0}$, and so the $P$–orbit closure of almost every point of $\mathcal{Q}\mathbb{T}^{0}$ is $\mathcal{L}$. ∎
Remark 4.4.
We note that the above proof also works for any subtorus $\mathcal{Q}\mathbb{T}^{\prime}\subset\mathcal{Q}\mathbb{T}^{0}$; the point is just that the intersection of any AIS $\mathcal{N}$ that does not contain $\mathcal{Q}\mathbb{T}^{\prime}$ with the torus is always a union of subtori. Thus the generic point of $\mathcal{Q}\mathbb{T}^{\prime}$ equidistributes under the $P$ action to $\overline{P\cdot\mathcal{Q}\mathbb{T}^{\prime}}$.
4.3. Individual pushforwards
Forni proved that whenever one has a nice $\mathsf{SL}_{2}\mathbb{R}$ action and equidistribution of the form above, one can remove the $T$-average in the above theorem at the cost of removing a small set of times. We recall that a set $Z\subset\mathbb{R}$ is said to have zero (upper) density if
$\limsup_{n\to\infty}\frac{\text{Leb}(Z\cap[0,n])}{n}=0$
and that a sequence $(a_{t})$ converges as $t\to\infty$ outside of $Z$ if it converges along all sequences tending to $\infty$ that are eventually disjoint from $Z$.
Theorem 4.5 (Theorem 1.1 of [For21]).
Let $\mathcal{Q}$ be a stratum of quadratic differentials and suppose that $\nu$ is a $U$-invariant probability measure on $\mathcal{Q}$. If the limit
[…]
exists and is $U$-ergodic, then there exists a set $Z\subset\mathbb{R}$ of zero density such that $(g_{t})_{*}\nu\to\nu_{\infty}$ as $t\to\infty$ outside of $Z$.
Remark 4.6.
Forni conjectures that $Z$ is empty in the case that $\nu_{\infty}$ is $P$-invariant.
Applying this to our context, we immediately get the following strengthening of Proposition 4.3:
Corollary 4.7.
With all notation as in Proposition 4.3, there is a set $Z\subset\mathbb{R}$ of zero density such that
$(g_{t})_{*}\nu_{\gamma}(\mathsf{Y},\mathbf{h})\to\nu_{\mathcal{L}}$
as $t\to\infty$ outside of $Z$.
Forni also proved that if $\nu$ is especially nice, then the set $Z$ is empty. Recall that a horospherical measure $\nu$ is one that is supported in a leaf of the unstable foliation of $g_{t}$ on a stratum $\mathcal{Q}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$, is absolutely continuous with respect to the canonical affine measure on the leaf, has continuous density, and has conditional measures along horocycle orbits equal to the restriction of one-dimensional Lebesgue measures.131313 Theorem 1.6 in Forni’s paper is stated for the stable foliation, but clearly applies to the unstable.
Proof of Theorem 1.5.
From the definitions and the fact that $\mathcal{O}$ is a piecewise real-analytic homeomorphism on unstable leaves that takes earthquake flow to horocycle flow in a time-preserving fashion (Theorems 2.1 and 2.3), we see that if $\sigma$ is a horospherical probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h})$ (in the sense of §3.3), then $\mathcal{O}_{*}\sigma$ is horospherical with support $\mathcal{Q}H_{\gamma}(\mathbf{h})$ (in the sense above). Similarly, if $\sigma$ is a horospherical slice probability measure supported on $\mathcal{P}H_{\gamma}(\mathbf{h},\Gamma)$, then $\mathcal{O}_{*}\sigma$ is supported on $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ and is horospherical for the ambient stratum component.
Theorem 1.6 of [For21] (or [EMM22, Proposition 3.2]) ensures $(g_{t}\circ\mathcal{O})_{*}\sigma$ converges weak-* to the Masur–Smillie–Veech measure $\nu_{\infty}$ on the ambient stratum as $t\to\infty$. Applying Theorem 1.4 then gives that $(\operatorname{Dil}_{t})_{*}\sigma$ converges weak-* to $\mathcal{O}^{*}\nu_{\mathcal{Q}}$, where $\mathcal{Q}$ is Masur–Smillie–Veech measure on the stratum component of $\mathcal{Q}^{1}\mathcal{M}_{g}$ containing the horosphere (slice). ∎
5. Identifying orbit closures
Our goal is now to identify $\mathcal{L}=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\underline{A},\mathbf{h})}$. For this, we will need the notion of the rank of an AIS. Let $\widehat{g}\geq 2$ and suppose ^$\widehat{}\underline{\kappa}$ is a partition of $22\widehat{g}-$ 2. We consider the unit area locus $\widehat{\mathcal{L}}$ of an AIS in the stratum (^$\mathcal{H}(\widehat{}\underline{\kappa})$ of abelian differentials determined by ^$\widehat{}\underline{\kappa}$. Then $\mathbb{R}\widehat{\mathcal{L}}$ is an immersed submanifold of (^$\mathcal{H}(\widehat{}\underline{\kappa})$ that is locally a $\mathbb{C}$-linear subspace in period coordinates for (^$\mathcal{H}(\widehat{}\underline{\kappa})$ defined over $\mathbb{R}$ (see §4.2). For $\omega\in\mathbb{R}\widehat{\mathcal{L}}$, we have
$T_{\omega}\mathbb{R}\widehat{\mathcal{L}}\subset T_{\omega}\mathcal{H}(\widehat{}\underline{\kappa})\cong H^{1}(\omega,Z(\omega);\mathbb{C}),$
and there is a natural linear projection $p$ from relative to absolute cohomology.
Definition 5.1.
The rank of $\widehat{\mathcal{L}}$ is the half the complex dimension of $p(T_{\omega}\mathbb{R}\widehat{\mathcal{L}})\subset H^{1}(\omega;\mathbb{C})$. An AIS is said to have high rank if its rank is at least $\widehat{g}/2+1$.
Recent deep work of Apisa and Wright classifies the AISs of high rank.
Theorem 5.2 (Theorem 1.1 of [AW23]).
Let $\widehat{\mathcal{L}}$ be an AIS in a stratum (^$\mathcal{H}(\widehat{}\underline{\kappa})$ of genus $\widehat{g}$ abelian differentials. Then if $\widehat{\mathcal{L}}$ has high rank, then either $\widehat{\mathcal{L}}$ is a component of (^$\mathcal{H}(\widehat{}\underline{\kappa})$ or the locus of all holonomy double covers of surfaces in a component of a stratum of quadratic differentials.
We would like to apply Theorem 5.2 to identify $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$, the $P$-orbit closure of a flat twist torus or horosphere slice. To do so, consider the locus ^(^$\widehat{\mathcal{L}}\subset\mathcal{H}(\widehat{}\underline{\kappa})$ of holonomy double covers, which is itself an AIS. Let $\underline{\kappa}_{\operatorname{odd}}$ denote those elements of $\underline{\kappa}$ that are odd.
If $\varepsilon=+1$ (hence #$\#\underline{\kappa}_{\operatorname{odd}}=0$), then $\widehat{q}\to q$ is the trivial double cover, $\widehat{\mathcal{Q}}\cong\mathcal{Q}$, and the rank of $\widehat{\mathcal{Q}}$ is $g$. Otherwise, we record the relevant numerics below.
Lemma 5.3.
Let $\mathcal{Q}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$ be a component of a stratum and $\widehat{\mathcal{Q}}\subset\mathcal{H}(\widehat{}\underline{\kappa})$ the locus of holonomy double branched covers. Then ^$\widehat{}\underline{\kappa}$ is a partition of $22\widehat{g}-$ 2, where
[…]
Proof.
This is [AW21, Lemma 4.2], but we give some details for completeness; see also [BW17, Theorem 26] for a related computation in the context of train tracks. The expression for $\widehat{g}$ can be computed using the Riemann–Hurwitz formula for the holonomy double cover $\widehat{q}\to q$ branched over the zeros of $q\in\mathcal{QM}(\underline{\kappa};-1)$. To compute rank, subtract the complex dimension of the kernel of the map to absolute cohomology from the complex dimension of $\mathcal{QM}(\underline{\kappa};-1)$ and divide by two. Since $\dim\mathcal{QM}(\underline{\kappa};-1)=2g-2+\#\underline{\kappa}$ and the kernel of the map to absolute cohomology is generated by certain relative cycles associated to the even order zeros of any $q\in\mathcal{Q}$, the formula follows. ∎
Corollary 5.4.
Suppose that $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$ is an AIS and that
$\text{rank}(\mathbb{R}\widehat{\mathcal{L}})\geq g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}.$
Then $\mathcal{L}$ is a component of $\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$.
Proof.
Clearly $\widehat{\mathcal{L}}$ is contained inside the locus of all holonomy double covers of differentials in $\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$. The rank hypothesis together with Lemma 5.3 implies that
$\text{rank}(\widehat{\mathcal{L}})\geq g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}=\widehat{g}/2+1,$
hence $\mathbb{R}\widehat{\mathcal{L}}$ has high rank and so we can apply Theorem 5.2. ∎
Remark 5.5.
We note that, by the second computation in Lemma 5.3, the hypotheses of the Corollary can only be satisfied when there are at least 6 odd order zeros.
5.1. Cylinder deformations and rank
To apply Theorem 5.2, we must bound the rank of $\mathcal{L}$ containing a twist torus or a horosphere slice from below. For this, we use cylinder deformations, which were introduced in [Wri15] and are recalled below.
For a given quadratic differential $q$, let $\mathcal{C}$ be the set of maximal, singularity free horizontal cylinders in $q$. Each element $C_{i}\in\mathcal{C}$ is an isometrically embedded open cylinder with height $h_{i}>0$ and circumference $\ell_{i}>0$, and $C_{i}$ is foliated by closed nonsingular leaves of the imaginary foliation of $q$ which are all homotopic to its core curve $\gamma_{i}$. Recall that
$u_{s}=\begin{pmatrix}1&s\\ 0&1\end{pmatrix}\in\operatorname{\mathsf{PSL}}_{2}\mathbb{R},$
and define $u_{s}^{C_{i}}q$ to be the holomorphic quadratic differential obtained by applying the horocycle flow $u_{s}$ to $C_{i}$ but not the rest of $q$.
For the rest of this section, $\Theta_{\gamma}\subset\mathcal{L}$ denotes either a (flat) twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ or a horosphere (slice) $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ as constructed in §4. Then each $q\in\Theta_{\gamma}$ is completely horizontally periodic, the core curves of the horizontal cylinders $\mathcal{C}$ constitute $\gamma=\cup_{i=1}^{n}\gamma_{i}$, and the height of the cylinder $C_{i}$ corresponding to $\gamma_{i}$ is $h_{i}$. From the definitions, we see that
$u_{s_{1}}^{C_{1}}\cdot...\cdot u_{s_{n}}^{C_{n}}q\in\Theta_{\gamma}$
for all $(s_{1},...,s_{n})\in\mathbb{R}^{n}$. Denote by $\eta_{i}$ the tangent vector to the path $s\mapsto u_{s}^{C_{i}}q$ at $s=0$, so that
(6) $\langle\eta_{1},...,\eta_{n}\rangle\leq T_{q}\Theta_{\gamma}\leq T_{q}\mathcal{L}\cong T_{\widehat{q}}\widehat{\mathcal{L}}.$
The linear holonomy along the core curve of a horizontal cylinder is trivial, so we know the preimage of $\gamma_{i}$ in the holonomy cover $\widehat{q}$ has two components. Pick a lift $\overline{\gamma}_{i}\subset\widehat{q}$ so that the full preimage is $\overline{\gamma}_{i}\sqcup\iota(\overline{\gamma}_{i})$, where $\iota$ denotes the covering involution. Choose an orientation on each $\gamma_{i}\subset S$ and orient $\overline{\gamma}_{i}$ so that it maps with degree $1$ to $\gamma_{i}$. Push that orientation forward by $\iota$ to induce an orientation on $\iota(\overline{\gamma}_{i})$ and define $\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}\in H_{1}(\widehat{q}\setminus Z(\widehat{q});\mathbb{Z})$. It satisfies $\iota_{*}\widehat{\gamma}_{i}=-\widehat{\gamma}_{i}$. See Figure 6.
The following Lemma is the content of [Wri15, §2] (after lifting the situation from $q$ to $\widehat{q}$).
Lemma 5.6.
Up to multiplication by ±$\pm 1$, the deformation $\eta_{i}\in H^{1}(\widehat{q},Z(\widehat{q});\mathbb{C})^{-}\cong T_{\widehat{q}}\mathbb{R}\widehat{\mathcal{Q}}$ is Poincaré–Lefschetz dual to the cylinder $h_{i}\widehat{\gamma}_{i}\in H_{1}(\widehat{q}\setminus Z(\widehat{q});\mathbb{C})^{-}$.
The inclusion induces a natural map $i_{*}:H_{1}(\widehat{q}\setminus Z(\widehat{q});\mathbb{Z})\to H_{1}(\widehat{q};\mathbb{Z})$ and Poincaré–Lefschetz duality makes the following square commute (with any coefficients):
[…]
Proposition 5.7.
The rank of any AIS $\mathcal{L}$ containing $\Theta_{\gamma}$ is bounded below by the real dimension of
$i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle_{\mathbb{R}}\leq H_{1}(\widehat{q};\mathbb{R})^{-}.$
Proof.
Lemma 5.6 together with (6) give that
(7) $p(\langle\eta_{1},...,\eta_{n}\rangle)=i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle_{\mathbb{R}}\otimes\mathbb{C}.$
By [AEM17], $p(T_{\widehat{q}}\widehat{\mathcal{L}})$ is symplectic, and since $\gamma_{i}\cap\gamma_{j}=\emptyset$ for $i\not=j$, the subspace (7) is isotropic. Thus the complex dimension of $p(T_{\widehat{q}}\widehat{\mathcal{L}})$ is bounded below by twice the complex dimension of (7). ∎
5.2. Lifting homologies
Proposition 5.7 reduces our problem to a topological question about (the lift of) $\gamma$ in the surface $\widehat{S}$ homeomorphic to $\widehat{q}$. The topology of $\widehat{S}$ depends not just on $\gamma$, but also on the geometry of the complement of $\gamma$ in the corresponding hyperbolic/flat structures. The relevant data is encoded in the geometric arc system $\underline{\alpha}$ dual to the ribbon graph $\mathsf{Y}$ or $\Gamma$.
Definition 5.8.
Let $\Sigma$ be a compact surface with non-empty boundary and let $\underline{\alpha}$ be a filling arc system on $\Sigma$. Say that $\underline{\alpha}$ is co-orientable on $\Sigma$ if there is an orientation of $\partial\Sigma$ and a co-orientation of $\underline{\alpha}$ continuously extending the orientation on $\partial\Sigma$.
Given a multicurve $\gamma\subset S$ and a filling arc system $\underline{\alpha}$ on $S\setminus\gamma$, the pair $(\gamma,\underline{\alpha})$ is jointly orientable if $\underline{\alpha}$ can be co-oriented on all complementary subsurfaces inducing a consistent orientation on $\gamma$.
Clearly, if $(\gamma,\underline{\alpha})$ is jointly orientable, then the restriction of $\underline{\alpha}$ to each component of $S\setminus\gamma$ is co-orientable. On the other hand, $(\gamma,\underline{\alpha})$ need not be jointly orientable even if the restriction of $\underline{\alpha}$ to each component of $S\setminus\gamma$ is co-orientable. See Figure 4.
Figure 4. A co-orientable $\underline{\alpha}$ on $S\setminus\gamma$ where $(\gamma,\underline{\alpha})$ is not jointly orientable.
Co-orientability and joint orientability are related to the topology of $\widehat{S}\to S$ as follows.
Lemma 5.9.
Let $q\in\Theta_{\gamma}$ and consider the corresponding holonomy double branched cover $\widehat{S}\to S$. For any component $\Sigma\subset S\setminus\gamma$, the preimage $\widehat{\Sigma}$ is disconnected if and only if $\underline{\alpha}|_{\Sigma}$ is co-orientable, and $\widehat{S}$ is disconnected if and only if $(\gamma,\underline{\alpha})$ is jointly orientable.
Proof.
We begin by observing that we can fix nice representatives for the components of $S\setminus\gamma$ and for $\underline{\alpha}$ as follows. By construction, each curve of $\gamma$ can be realized as a core curve of a horizontal cylinder of $q$. Cutting along these core curves results in a flat cone surface with boundary homeomorphic to $S\setminus\gamma$. Each component $\Sigma$ obtained this way deformation retracts onto the graph of horizontal saddle connections contained in $\Sigma$ by collapsing the segments of its vertical foliation, and we can pick a representative of $\underline{\alpha}$ by choosing one segment in each isotopy class. Compare Figure 5.
From this description, it follows that $\underline{\alpha}$ is co-orientable on $\Sigma$ if and only if the restriction of the horizontal foliation $\operatorname{Im}(q)$ to $\Sigma$ is orientable (see also the discussion in Section 7.2 of [CF24b]), which happens if and only if the preimage $\widehat{\Sigma}$ is disconnected. Similarly, a joint orientation on $(\gamma,\underline{\alpha})$ is equivalent to a global choice of orientation of $\operatorname{Im}(q)$ on $S$, which implies the holonomy double cover is disconnected. ∎
Figure 5. A co-orientation on $\underline{\alpha}|_{\Sigma}$ and the corresponding orientation of the horizontal foliation of $q$ restricted to $\Sigma$.
The following Proposition is proved by a technical homology computation in $\widehat{S}$. The basic idea is that each subsurface of $S\setminus\gamma$ defines a linear relation between its boundary curves in homology on $S$, and these relations lift to relations between the curves of $\widehat{\gamma}$ in $\widehat{S}$ exactly when that subsurface is disconnected in $\widehat{S}$. The main difficulty in the proof comes from ensuring that there are no other unaccounted-for relations.
Proposition 5.10.
Any invariant subvariety $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$ containing $\Theta_{\gamma}$ with defining arc system $\underline{\alpha}$ (dual to $\mathsf{Y}$ or $\Gamma$) has rank at least
#$\#\gamma-N_{co}+\delta_{jo},$
where $N_{co}$ is the number of components $\Sigma$ of $S\setminus\gamma$ for which $\underline{\alpha}|_{\Sigma}$ is co-orientable, and $\delta_{jo}=1$ if $(\gamma,\underline{\alpha})$ is jointly orientable and 00 otherwise.
Proof.
We will show that #$\#\gamma-N_{co}+\delta_{jo}$ is precisely the dimension of the space
$V=i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle\leq H_{1}(\widehat{S};\mathbb{R})^{-},$
so that Proposition 5.7 establishes the relationship with rank. For brevity, and to avoid confusion with the action $\iota_{*}$ of the covering involution on the homology of $\widehat{S}$, we drop the $i_{*}$ and write $\widehat{\gamma}_{i}\in H_{1}(\widehat{S};\mathbb{R})^{-}$.
Enumerate the subsurfaces $S\setminus\gamma=\Sigma_{1}\cup...\cup\Sigma_{k}$, and let $\widehat{\Sigma}_{i}$ be the full preimage of $\Sigma_{i}$ under $\widehat{S}\to S$. Consider annular neighborhoods $A_{i}$ of $\gamma_{i}\subset S$ and let $\widehat{A}_{i}$ be the full preimages. Recall that the anti-invariant part of $H_{1}(\widehat{A}_{i})$ is generated by $\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}$.
Figure 6. A non-trivial cover $\widehat{\Sigma}\to\Sigma$ corresponding to an arc system $\underline{\alpha}|_{\Sigma}$ that is not co-orientable. We have $\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}$ and the push-off to the right is $\widehat{\gamma}_{i}^{+}=\overline{\gamma}_{i}^{+}-\iota_{*}\overline{\gamma}_{i}^{+}$.
By Lemma 5.9, a subsurface $\widehat{\Sigma}_{j}$ is disconnected if and only if $\underline{\alpha}|_{\Sigma_{j}}$ is co-orientable. Our claim is then that the relation among the boundary curves of $\Sigma_{j}$ in homology yields a relation among the corresponding homology classes in $\widehat{\Sigma}_{i}$ when $\underline{\alpha}|_{\Sigma_{j}}$ is co-orientable, and otherwise the relation does not lift. We also claim that these are the only relations satisfied by the $\{\widehat{\gamma}_{i}\}$, and that there is even a redundancy among them in the case that $(\gamma,\underline{\alpha})$ is jointly orientable.
To make this precise, consider the Mayer–Vietoris long exact sequence (with coefficients in $\mathbb{R}$, say)
(8) $0\to H_{2}(\widehat{S})\to\oplus_{i,j}H_{1}(\widehat{A}_{i}\cap\widehat{\Sigma}_{j})\to\left(\oplus_{i=1}^{n}H_{1}(\widehat{A}_{i})\right)\bigoplus\left(\oplus_{j=1}^{k}H_{1}(\widehat{\Sigma}_{j})\right)\to H_{1}(\widehat{S})\to...$
By naturality of $\iota_{*}$, this long exact sequence is the sum of two long exact sequences corresponding to the $\iota$-invariants and the $\iota$-anti-invariants, and we are interested in the latter.
Assume first that $\widehat{S}$ is connected, i.e., that $(\gamma,\underline{\alpha})$ is not jointly orientable (Lemma 5.9). Since $\iota$ is orientation-preserving and $\widehat{S}$ is closed, we have $H_{2}(\widehat{S})^{-}=0$. From the definitions, $H_{1}(\widehat{A}_{i})^{-}=\langle\widehat{\gamma}_{i}\rangle$. Let $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ be curves parallel to $\gamma_{i}$ obtained by pushing off to the right and left, respectively. For any fixed $i$, we have that $\bigcup_{j}\widehat{A}_{i}\cap\widehat{\Sigma}_{j}$ is a union of four cylinders, the cores of which correspond to $\overline{\gamma}_{i}^{\pm}$ and $\iota_{*}\overline{\gamma}_{i}^{\pm}$. Thus $\oplus_{i,j}H_{1}(\widehat{A}_{i}\cap\widehat{\Sigma}_{j})^{-}$ is freely generated by $\{\widehat{\gamma}_{i}^{\pm}:i=1,...,n\}$.
Recall that $\widehat{\Sigma}$ is the total preimage of $\Sigma$ under the two-fold branch covering projection, which is then invariant under the covering involution. Hence the anti-invariance distributes over the sum in (8), yielding
[…]
where $F:\widehat{\gamma}_{i}^{\pm}\mapsto\widehat{\gamma}_{i}-\widehat{\gamma}_{i}^{\pm}$ and $G$ sums the components together. Note $\dim\operatorname{im}F=2n$ since (9) is exact.
We now use this sequence to compute the dimension of $V$. We begin by observing that $\operatorname{im}F\subset G^{-1}(V)$, so that $V\cong G^{-1}(V)/\operatorname{im}F$. From the definition of $V$, we see $G^{-1}(V)$ is generated by the $3n$ homology classes $\{\widehat{\gamma}_{i}^{-},\widehat{\gamma}_{i},\widehat{\gamma}_{i}^{+}\}_{i=1}^{n}$. The classes $\{\widehat{\gamma}_{i}^{\pm}\}_{i=1}^{n}$ may satisfy some relations in $\oplus H_{1}(\widehat{\Sigma}_{j})^{-}$, which we now compute.
Each $\widehat{\Sigma}_{j}$ is a surface with an even number $m_{j}$ of boundary components
$\{c_{1},...,c_{m_{j}}\}\cup\{\iota_{*}c_{1},...,\iota_{*}c_{m_{j}}\},$
where $\{c_{i}\}\subset\{\pm\overline{\gamma}_{i}^{\pm}\}$. We assume without loss of generality that the orientations are chosen so that
$\sum c_{i}+\sum\iota_{*}c_{i}=0\in H_{1}(\widehat{\Sigma}_{j})$
holds. If $\widehat{\Sigma}_{j}$ is connected, then this is the only non-trivial relation satisfied by $\{c_{i}\}\cup\{\iota_{*}c_{i}\}$, while if $\Sigma_{j}$ is disconnected, then additionally
$\sum c_{i}=0=\sum\iota_{*}c_{i}\in H_{1}(\widehat{\Sigma}_{j}).$
In particular,
$\{c_{i}-\iota_{*}c_{i}:{i=1,...,m_{j}}\}\subset H_{1}(\widehat{\Sigma}_{j})^{-}$
is a linearly independent set if and only if $\widehat{\Sigma}_{j}$ is connected and otherwise spans a subspace of dimension $m_{j}-1$. By definition, there are $N_{co}$ such disconnected $\widehat{\Sigma}_{j}$, so that $\dim G^{-1}(V)=3n-N_{co}$.
Since $\widehat{S}$ is connected by assumption, i.e., $(\gamma,\underline{\alpha})$ is not jointly orientable and $\delta_{jo}=0$, we have that $V\cong G^{-1}(V)/\operatorname{im}F$ has dimension
$\dim G^{-1}(V)-\dim\operatorname{im}F=(3n-N_{co})-2n=n-N_{co}+0=\#\gamma-N_{co}+\delta_{jo}.$
In the case that $\widehat{S}$ is disconnected, i.e., $\delta_{jo}=1$, the computation is identical except that $H_{2}(\widehat{S})^{-}\cong H_{2}(S)$, and so $\operatorname{im}F$ has dimension $12n-$ 1. Thus $V$ has dimension $3n-N_{co}-(2n-1)=\#\gamma-N_{co}+\delta_{jo}$, completing the proof of the Proposition. ∎
5.3. Orbit closures of twist tori
We now apply the computation of the previous subsection. Let $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ be a flat twist torus constructed from a multicurve $\gamma$ (not necessarily a pants decomposition). Let $\underline{\alpha}$ be the arc system dual to $\mathsf{Y}$, and let $\underline{\kappa}$ be the partition of $44g-$ 4 determined by the valences of the vertices of $\mathsf{Y}$. Let $\varepsilon=1$ if $(\gamma,\underline{\alpha})$ is jointly orientable and $\varepsilon=-1$, otherwise.
Corollary 5.11.
Let $\mathcal{L}=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})}$. If $\varepsilon=-1$ and
$g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}\leq\#\gamma-N_{co},$
then $\mathcal{L}$ is the ambient component of $\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$ containing $\mathcal{L}$.
Proof.
That $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$, hence that $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$, follows from the constructions and the definitions.
Let $V=i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle\leq H_{1}(\widehat{S};\mathbb{R})$ be as in Proposition 5.7 so that $\text{rank}(\mathcal{L})\geq\dim_{\mathbb{R}}V$ holds. Proposition 5.10 computes $\dim_{\mathbb{R}}V=\#\gamma-N_{co}$ when $(\gamma,\underline{\alpha})$ is not jointly orientable (i.e., when $\varepsilon=-1$), and Lemma 5.3 computes $\widehat{g}/2+1=g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}$. Corollary 5.4 applies and we conclude that $\mathcal{L}=\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$. ∎
We now specialize this to the setting of Conjecture 1.1. Given a pants decomposition $\gamma$ of a topological surface $S$ and a length vector $\bm{\ell}\in\mathbb{R}_{>0}^{3g-3}$, let $\nabla_{\gamma}(\bm{\ell})$ denote the number of components of $S\setminus\gamma$ that have property ()($\nabla$)( ∇ ) from the introduction, i.e., such that the lengths of two of its boundaries sum to the length of the third (this exactly corresponds to the spine having a 4-valent vertex and $\underline{\alpha}$ being co-orientable on that component).
Corollary 5.12.
For any pants decomposition $\gamma$, any $\bm{\ell}\in\mathbb{R}_{>0}^{3g-3}$, and any $\mathbf{h}\in\mathbb{R}_{>0}^{3g-3}$ satisfying $\bm{\ell}\cdot\mathbf{h}=1$, set
$\mathcal{L}:=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})}.$
  1. (1)
    If $\nabla_{\gamma}(\bm{\ell})\leq 2g-5$, then $\mathcal{L}$ is the (connected) stratum $\mathcal{Q}^{1}\mathcal{M}_{g}(1^{4g-4-2\nabla_{\gamma}(\bm{\ell})},2^{\nabla_{\gamma}(\bm{\ell})};-1)$.
  2. (2)
    If $\nabla_{\gamma}(\bm{\ell})=2g-2$ and if $(\gamma,\underline{\alpha}(X))$ is jointly orientable for some (any) $X\in\mathbb{T}_{\gamma}(\bm{\ell})$, then $\mathcal{L}$ is a component of $\mathcal{Q}^{1}\mathcal{M}_{g}(2^{2g-2};+1)$.
Proof.
The first statement follows immediately from Corollary 5.11. For the second, we note that in this case the corresponding twist torus lives in a stratum consisting of squares of abelian differentials. Since each component of $S\setminus\gamma$ has genus 00, the curves of $\gamma$ span a $g$-dimensional subspace of $H_{1}(S)$, hence $\mathcal{L}$ has rank $g$. Work of Mirzakhani and Wright [MW17] classifies rank $g$ AISs in strata of abelian differentials as either the entire stratum component or the hyperelliptic locus. One then checks by inspection that if a twist torus lives entirely in the hyperelliptic locus then each of its curves must be fixed by the hyperelliptic involution, which is impossible when $\gamma$ is a pants decomposition. ∎
We now assemble the proof of (a more detailed version of) our main theorem.
Proof of Theorem 1.2.
Let $\gamma$ be any pants decomposition and fix any $\bm{\ell}$ and $\mathbf{h}$ as in the statement of the Theorem. Since pants are determined by their boundaries, $\bm{\ell}$ specifies a hyperbolic structure on $S\setminus\gamma$; let $\mathsf{Y}$ denote its spine and $\underline{\alpha}$ the dual arc system. We note that the edge lengths of $\mathsf{Y}$ are a linear function of the lengths $\bm{\ell}$ (see [CF24b, Figure 1]), so it is equivalent to rescale $\bm{\ell}$ and to rescale $\mathsf{Y}$.
Form the twist tori $\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ and $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ as in §§3.1 and 4.1, as well as their associated measures $\mu_{\gamma}(\mathsf{Y},\mathbf{h})$ and $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$. Recall that $\mathcal{O}$ interchanges these tori and that $g_{t}$ takes twist tori to twist tori (5):
$\mathcal{O}_{*}\mu_{\gamma}(\mathsf{Y},\mathbf{h})=\nu_{\gamma}(\mathsf{Y},\mathbf{h})\hskip 8.61108pt\text{and}\hskip 8.61108pt(g_{t})_{*}\nu_{\gamma}(\mathsf{Y},\mathbf{h})=\nu_{\gamma}(e^{t}\mathsf{Y},e^{-t}\mathbf{h}).$
The spine $\mathsf{Y}$ has $22g-$ 2 components, $\nabla_{\gamma}(\bm{\ell})$ of which have a unique 4-valent vertex and the rest of which have two trivalent vertices. By Theorem 2.3 (3), we can read off the stratum containing the flat twist torus:
$\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})\subset\mathcal{Q}^{1}\mathcal{M}_{g}(1^{4g-4-2\nabla_{\gamma}(\bm{\ell})},2^{\nabla_{\gamma}(\bm{\ell})};\varepsilon)$
where $\varepsilon=+1$ if $(\gamma,\underline{\alpha})$ is jointly orientable and $-1$ otherwise.
By Corollary 4.7 and (5), we have that these flat twist tori equidistribute outside a set $Z$ of zero density:
[…]
where $\mathcal{L}$ is the $P$-orbit closure of $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. Applying Theorem 1.4, we therefore get that
[…]
Since $\nu_{\mathcal{L}}$ lives in the same ambient stratum as $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$, this proves part (2) of Theorem 1.2.
Of course, we can do even better and identify the orbit closure in most cases. Corollary 5.12 tells us that if $\nabla_{\gamma}(\bm{\ell})\leq 2g-5$, or $\nabla_{\gamma}(\bm{\ell})=2g-2$ and we have joint orientability, then $\mathcal{L}$ is the entire ambient stratum component. In particular, if $\nabla_{\gamma}(\bm{\ell})=0$ we have that $\mathcal{L}$ is the principal stratum, and so
[…]
where the last equality comes from Theorem 2.3 (2). This completes the proof of the Theorem. ∎
5.4. Other high rank twist tori
Here are some more concrete families of hyperbolic surfaces where the same argument applies. The proofs of the following results are identical to the proof of Theorem 1.2 (apart from some arithmetic to check that one can apply Corollary 5.11) and are therefore left to the reader.
In the following examples, the limiting distributions on $\mathcal{P}^{1}\mathcal{M}_{g}$ are qualitatively different from Mirzakhani measure. For example, the generic surface can contain large embedded disks, can be hyperelliptic, can have constrained thick-thin decomposition, and so on. We have chosen these families to be illustrative, not exhaustive — the reader is encouraged to come up with their own.
Regular plumbing fixtures
For $p\geq 3$, a regular plumbing fixture is a hyperbolic metric on a $p$-holed sphere with totally geodesic boundary and a dihedral symmetry group of order $2p$; see Figure 7. Regular plumbing fixtures with $p$ holes are parameterized by the length $L$ of each of their boundaries.
Figure 7. A regular plumbing fixture on a $7$-holed sphere.
Suppose that $\gamma$ is a multicurve on $S$ such that each component $\Sigma_{1},\ldots,\Sigma_{k}$ of $S\setminus\gamma$ has genus 0; let $p_{i}$ denote the number of boundary components of $\Sigma_{i}$. Taking each component to be a regular plumbing fixture with boundary lengths $L$, we get a twist torus whose corresponding metric ribbon graph $\mathsf{X}=\mathsf{X}(p_{1},\ldots,p_{k})$ has $k$ components, the $i^{\text{th}}$ of which has two vertices of valence $p_{i}$ and edges of length $L/2$.
Theorem 5.13.
Let all notation be as in the previous paragraphs and fix any $\mathbf{h}\in\mathbb{R}_{>0}^{\#\gamma}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. Then if at least 3 of the $p_{i}$ are odd, the twist tori $\mathcal{P}\mathbb{T}_{\gamma}(e^{t}\mathsf{X},e^{-t}\mathbf{h})$ equidistribute in $\mathcal{P}^{1}\mathcal{M}_{g}$ outside of set $Z$ of zero density to the pullback of Masur–Smillie–Veech measure on the corresponding stratum. That is,
[…]
where $\nu_{\mathcal{Q}}$ is the Masur–Smillie–Veech measure on $\mathcal{Q}=\mathcal{Q}^{1}\mathcal{M}_{g}\left((p_{1}-2)^{2},\ldots,(p_{k}-2)^{2};-1\right)$.
We note that if at least $3$ of the $p_{i}$ are odd (which is what allows us to apply Corollary 5.4), then by the classification of strata of quadratic differentials [Lan08, CM14] we know that $\mathcal{Q}$ must be connected.
As in Corollary 5.12, we can also deduce a complementary result when the twist torus lands in a stratum of squares of abelian differentials. However, in this case the gluing pattern determines the limiting measure.
Theorem 5.14.
If all $p_{i}$ are even and $(\gamma,\underline{\alpha})$ is jointly orientable, where $\underline{\alpha}$ is dual to $\mathsf{X}$, then
[…]
If $S$ admits an involution fixing each of the curves of $\gamma$ setwise but reversing their orientations (hence $k\leq 2$), then $\mathcal{L}$ is the hyperelliptic locus in the component of the stratum $\mathcal{Q}^{1}\mathcal{M}_{g}\left((p_{1}-2)^{2},\ldots,(p_{k}-2)^{2};+1\right)$ that contains $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{X},\mathbf{h})$. Otherwise, $\mathcal{L}$ is the entire component.
Remark 5.15.
If $S$ admits a hyperelliptic involution $\iota$ that fixes $\gamma$ as a multicurve (but not necessarily each of its components), then one can consider the maximal $\iota$-invariant subtorus inside of $\mathbb{T}_{\gamma}(e^{t}\mathsf{X})$. The same reasoning as above (combined with Remark 4.4) shows that these invariant subtori also equidistribute to the measure supported on the hyperelliptic locus in the ambient stratum.
Minimal strata
We can also construct flat twist tori whose $P$-orbit closures are different components of the minimal stratum of (squares of) abelian differentials. For example, let $\gamma$ be one of the multicurves shown in Figure 8. Its complement is a $2g$-holed sphere, and setting the weight of each blue arc equal to 1 determines a hyperbolic structure on $S\setminus\gamma$ all of whose boundaries have length $2$ except for one which has length $1$. The dual ribbon graph $\mathsf{X}$ has a unique vertex of valence $4g$ and all edges of length $1$. Moreover, it is easy to check by hand that the pair $(\gamma,\underline{\alpha})$ is jointly orientable and one can even read off the parity of the induced spin structure from the data of $\gamma$ and $\underline{\alpha}$ (see [CS21, §6.3]).
Figure 8. Twist tori corresponding to the minimal stratum of squares of abelian differentials. The leftmost torus lands in the hyperelliptic locus, while the other two land in the odd and even spin components (which diagram corresponds to which spin depends on the parity of $g$ mod $4$).
Theorem 5.16.
With $\gamma$ and $\mathsf{X}$ as in Figure 8 and any choice of $\mathbf{h}\in\mathbb{R}_{>0}^{\#\gamma}$ such that $\bm{\ell}\cdot\mathbf{h}=1$,
[…]
where $\mathcal{Q}$ is the component of the stratum $\mathcal{Q}^{1}\mathcal{M}_{g}(4g-4;+1)$ specified by the gluing pattern of $\gamma$.